### v2

# Involutive (simple) latin solutions of the Yang-Baxter equation and related (left) quasigroups

Abstract
In this paper, we study involutive non-degenerate set-theoretic solutions of the Yang-Baxter equation with regular displacement group. In particular, we completely describe the blocks of imprimitivity and the congruences of the irretractable ones, that we show belonging to the class of the latin solutions. Among these solutions, we characterise the simple ones having nilpotent permutation group. A more precise description involving the First Weyl Algebra will be provided when the displacement group is abelian and normal in the total permutation group, and we enumerate and classify the simple ones having minimal size $p^{p}$, for an arbitrary prime number $p$. Finally, we illustrate our results by some examples.
keywords:
imprimitive group , quasigroup, Yang-Baxter equation, brace, cycle set
MSC:
[2020] 16T25, 81R50 , 20N05
mytitlenotemytitlenotefootnotetext: The second author is a member of GNSAGA (INdAM).
Introduction
The quantum Yang-Baxter equation first appeared in theoretical physics, in a paper by C.N. Yang [44], and in statistical mechanics, in R.J. Baxter’s work [5]. On the other hand, it has been studied also been studied from a mathematical viewpoint. In 1992 Drinfel’d [24] suggested the study of set-theoretical solutions, i.e. solutions that act on a basis of the underlying vector space. Specifically, a set-theoretic solution of the Yang-Baxter equation on a non-empty set $X$ is a pair $\left(X,r\right)$, where $r:X\times X\to X\times X$ is a map such that the relation
$\left(r\times\operatorname{id}_{X}\right)\left(\operatorname{id}_{X}\times r\right)\left(r\times\operatorname{id}_{X}\right)=\left(\operatorname{id}_{X}\times r\right)\left(r\times\operatorname{id}_{X}\right)\left(\operatorname{id}_{X}\times r\right)$
is satisfied. Set theoretical solutions provide solutions of the Yang-Baxter equation by linearization. Writing a solution $(X,r)$ as $r\left(x,y\right)=\left(\lambda_{x}\left(y\right)\rho_{y}\left(x\right)\right)$, with $\lambda_{x},\rho_{x}$ maps from $X$ into itself, for every $x\in X$, we say that $(X,r)$ is non-degenerate if $\lambda_{x},\rho_{x}\in\operatorname{Sym}_{X}$, for every $x\in X$, and involutive if $r^{2}=\operatorname{id}_{X\times X}$.
The seminal papers due by Gateva-Ivanova and Van Den Bergh [26] and Etingov, Schedler, and Soloviev [25] encouraged several authors to the study of the involutive non-degenerate set-theoretic solutions (which we simply call involutive solutions). Recall that a solution $(X,r)$ is said to be decomposable if $X$ can be expressed as a disjoint union of two non-empty subsets $Y_{1}$ and $Y_{2}$ such that $r(Y_{i}\times Y_{i})\subseteq Y_{i}\times Y_{i}$ for all $i\in\{1,2\}$, and $(X,r)$ is called indecomposable, otherwise. To attack the classification-problem of involutive solutions, as first attempt one can restrict the investigations to the class of indecomposable ones. This approach is motivated by the fact that indecomposable solutions are in some sense the fundamental blocks that allow to understand all the involutive ones (see [25, Section $3$]). Even if we focus on indecomposable involutive solutions, a full classification is still hard, and in the last few years several authors investigated the so-called simple indecomposable involutive solutions, where a solution $(X,r)$ is said to be simple if every epimorphic image $(Y,s)$ is isomorphic to $(X,r)$ or is a singleton. A notion of simple involutive solution (not necessarily indecomposable) was at first introduced by Vendramin in [42] for involutive solutions not necessarily indecomposable. Even if his definition of simplicity is different, it coincides in the indecomposable case. Every simple involutive solution of size bigger than $2$ is indecomposable by [17]. These solutions play a special role since they are the “fundamental blocks” to construct all the indecomposable involutive ones by dynamical extensions (see [12, Proposition 2] and [42, Corollary 2.13]).
Even if the construction of simple involutive solutions appeared quite hard, several interesting results were obtained in the last few years. If $p$ is an odd prime number there exist a unique simple involutive solution of size $p$, see [25, Theorem 2.13]. For every natural number $n>1$, a simple involutive solution of cardinality $p^{n}$ was recently constructed in [19], and in the case $n=2$ a full classification was given in [22]. Further examples of cardinality $n^{2}$ and $mn^{2}$, where $n,m$ are arbitrary natural number, were constructed in [17, 20]. On the other hand, we know that an indecomposable involutive solutions of square-free cardinality can not be simple, as showed in [18].
In order to provide a theoretic description of simple solutions, in [10] the second author characterised them in terms of skew left braces provided by their associated permutation groups. Recall that a skew left brace, introduced in [27, 35], is a triple $(B,+,\circ)$ where $(B,+)$ and $(B,\circ)$ are groups and the equality
$x\circ(y+z)=x\circ y-x+x\circ z$
holds for all $x,y,z\in B$. The notion of simplicity was recently investigated in the general setting of bijective non-degenerate solutions (simply called solution) not necessarily involutive. In [13], a criterion of simplicity was provided by blocks of imprimitivity of the associated permutation group. A remarkable generalization of [10, Theorem $12$] was recently provided in [21], where the simple bijective solutions (not necessarily involutive) were completely characterised by the underlying permutation skew left braces, and several examples of non-involutive simple solutions were provided. In this way, the main Theorem of [21, Section $1$] represents a unifying tool to understand the notion of simplicity of several algebraic structures already present in literature. By these last results, it is clear that for the simplicity of a solution the ideal $B^{(2)}$ of the permutation left brace plays a fundamental role. However, we are still far away from a concrete description of the maps $r$ related to simple solutions. Indeed, in [21] a complete description of simple solutions provided by skew left braces $B$ with $B^{(2)}=0$ or $B^{(3)}=0$ was given, but the remaining case $B^{(2)}=B^{(3)}\neq 0$ is still open. As noted in [21, Section $1$], the last case includes the involutive simple solutions that are not of Lyuabashenko type. To partially fill this gap, one of the main goals of this paper is the study of finite involutive simple solutions in which $B^{(2)}$, that in this case coincides with the so-called displacement group, acts regularly on the underlying set. We will show that these solutions belong to the family of latin solutions, i.e. solutions in which the maps $\lambda_{x}$ provide a quasigroup structure on the set $X$ (see [9]).However, our investigation is not limited to simple solutions. Motivated by [16, Theorem $3.1$], where involutive solutions with primitive permutation group were completely classified, we focus on blocks of imprimitivity of involutive solutions (not necessarily simple). In addition to the main result of [16], which ensures the existence of a complete blocks system, we will show that a complete blocks system can be always obtained by the orbits of a minimal normal subgroup of the permutation group. To develop a theory useful to simple involutive solutions, in the second part we specialize to the irretractable ones. By tools that come from quasigroups theory, we decribe all the complete blocks system of irretractable involutive solutions with regular displacement group. Among these complete blocks systems, we distinguish the ones that give rise to congruences, providing an extension of a result involving the family of finite affine quasigroups (see [9, Lemma 4.8]) to a larger class of finite quasigroups. In order to understand more about congruences, using tools that come from the left quasigroups theory provided in [7], we develop a covering theory of left quasigroups that allows to recover the one given in [41] for indecomposable involutive solutions as a special case. This is the reason for which, throught the paper, we will use the language of cycle sets. Recall that a cycle set is a left quasigroup $(X,\cdotp)$ in which the equality $(x\cdotp y)\cdotp(x\cdotp z)=(y\cdotp x)\cdotp(y\cdotp z)$ holds for all $x,y,z\in X$. Cycle sets are useful to detect solutions of the Yang-Baxter equation: indeed, non-degenerate cycle sets bijectively correspond to non-degenerate involutive solutions, and by [34, Propositions 1-2] the correspondence is given by
$(X,\cdotp)\longrightarrow(X,r)\quad with\quad r(x,y):=(\sigma_{x}^{-1}(y),\sigma_{x}^{-1}(y)\cdot x).$ (1)
As one can expect, indecomposable (resp. simple) involutive solutions correspond to indecomposable (resp. simple) cycle sets. In the core of the paper, we focus on simple involutive solutions $(X,r)$ with nilpotent regular displacement group. We will show that, in this setting, $X$ must have a prime-power cardinality $p^{n}$ and the $p$-Sylow subgroup of the permutation group $\mathcal{G}(X)$ is normal. In this case we can reduce the classification to solutions for which $\mathcal{G}(X)$ is a $p$-group. A complete description of these solutions will be provided, while the ones having cyclic displacement group will be concretely classified. These results will be refined if, in addition, the displacement group is abelian and normal in the so-called total permutation group. In this last case, we exhibit a connection with irreducible representations of the First Weyl Algebra over a field with a prime number of elements. We remark that the First Weyl Algebra has already appeared in the study of involutive solutions, see [9, 39]. These solutions has always size $p^{k}$, for a prime number $p$ and a natural number $k$ divisible by $p$. Among these solutions, the ones having size $p^{p}$ will be completely classified. Finally, we illustrate our main results with some examples, and we use some of them to answer in the affirmative to [17, Question 7.6].
1 Left quasigroups
1.1 Basics
A left quasigroup is an algebraic structure given by a non-empty set $X$ together with a binary operation $\cdot:X\times X\longrightarrow X$, such that the map
$\sigma_{x}:y\mapsto x\cdot y$
[…]
The associated permutation group (also called left multiplication group) is defined as
$\mathcal{G}(X)=\langle\sigma_{x},\,x\in X\rangle\leq Sym(X).$
If such a group is transitive on $X$ we say that $X$ is indecomposable (the same property is also known as connectedness).
The map
$\delta_{x}:y\mapsto y\cdot x$
[…]
$\mathcal{TG}(X)=\langle\sigma_{x},\,\delta_{x},\,x\in X\rangle.$
An equivalence relation on $X$ is said to be a congruence if for every $x,y,x^{\prime},y^{\prime}\in X$ we have that if $x\sim x^{\prime}$ and $y\sim y^{\prime}$ then $x\cdotp y\sim x^{\prime}\cdotp y^{\prime}$ and $\sigma_{x}^{-1}(y)\sim\sigma_{x^{\prime}}^{-1}(y^{\prime})$. Accordingly the operation $[x]\cdot[y]=[x\cdot y]$ is well-defined and the quotient set $X/\alpha$ together with this operation is a left quasigroup.
Let $X,Y$ be left quasigroups. A homomorphism from $X$ to $Y$ is a map $p:X\rightarrow Y$ such that $p(x\cdotp y)=p(x)\cdotp p(y)$ for all $x,y\in X$. If $p$ is surjective, then it will be called epimorphism, while a bijective homomorphism is said to be isomorphism. An isomorphism from a left quasigroup $X$ to itself will be called automorphism and the automorphism group will be indicated by $Aut(X,\cdotp)$.
Epimorphisms and congruence are essentially the same thing. Indeed if $\alpha$ is a congruence, the canonical map $X\longrightarrow X/\alpha$ is an epimorphism. On the other hand, $\ker{p}=\{(x,y)\in X^{2}\,:\,p(x)=p(y)\}$ is a congruence of $X$ and $Y\cong X/\ker{p}$.
Lemma 1.
[1, Lemma 1.8] If $X,Y$ are left quasigroups and $p:X\rightarrow Y$ is a left quasigroups epimorphism, the assignment $\sigma_{x}\mapsto\sigma_{p(x)}$ extends to a groups epimorphism $\bar{p}:\mathcal{G}(X)\rightarrow\mathcal{G}(Y)$.
We can introduce a particular family of normal subgroups of the associated permutation group.
Definition 1.
[7] Let $X$ be a left quasigroup and $H$ be a normal subgroup of $\mathcal{G}(X)$. Then, $H$ is admissible if $\sigma_{x}^{-1}\sigma_{h(x)}\in H$ for all $x\in X,$ $h\in H$.
The lattice of admissible subgroups is a sublattice of the lattice of the normal subgroups of $\mathcal{G}(X)$. Note that, if $H$ is admissible, then $H$ is normal and so we have that $\sigma_{x}^{-1}\sigma_{h(x)}\in H$ for all $x\in X,$ $h\in H$ if and only if $\sigma_{x}\sigma^{-1}_{h(x)}\in H$ for all $x\in X,$ $h\in H$
According to [7] we can define a Galois connection between the congruence lattice of a left quasigroup and the sublattice of the admissible subgroups. Given a congruence $\alpha$ we can define
$\mathcal{G}^{\alpha}=\{h\in\mathcal{G}(X)\,:\,h(x)\,\alpha\,x\,\text{ for every }x\in X\}.$
Note that $\mathcal{G}^{\alpha}$ is nothing but the kernel of the induced group homomorphism from $\mathcal{G}(X)$ onto $\mathcal{G}(X/\alpha)$ defined in Lemma 1 and it is admissible according to [7, Corollary 1.9]. On the other hand, given an admissible subgroup $H$, we can define the equivalence relation
$\mathcal{O}_{H}=\{(x,h(x))\,:\,h\in H,\,x\in Q\},$
namely the equivalence relation with blocks given by the orbits of $H$. Such equivalence is a congruence [7, Lemma 1.8]. We usually denote the factor left quasigroup with respect to $\mathcal{O}_{H}$ just by $X/H$.
Moreover, given a congruence $\alpha$ we can define the subgroup $\mathrm{Dis}_{\alpha}$ as the normal closure of $\{\sigma_{x}\sigma_{y}^{-1}\,:\,x\,\alpha\,y\}$ in the group $\mathcal{G}(X)$. Such subgroups are admissible. We denote $\mathrm{Dis}_{1_{X}}$ just by $\mathrm{Dis}(X)$ and we refer to it as the displacement group of $X$. By [7, Section $1$], the quotient $\mathcal{G}(X)/\mathrm{Dis}(X)$ is cyclic, and in particular $\mathcal{G}(X)=\mathrm{Dis}(X)\langle\sigma_{x}\rangle$, for every $x\in X$.
We can define the binary relation $\sim_{\sigma}$ on a left quasigroup $X$ given by
$x\sim_{\sigma}y:\Longleftrightarrow\sigma_{x}=\sigma_{y}$
for all $x,y\in X$. The relation $\sim_{\sigma}$ is not a congruence in general, A left quasigroup $X$ is called irretractable if $\sim_{\sigma}=0_{X}$ (sometimes in the literature the left quasigroups with the same property are called faithful). A congruence $\alpha$ is contained in the relation $\sim_{\sigma}$ if and only if $\mathrm{Dis}_{\alpha}$ is the trivial group.
Indecomposable left quasigroups can always be described in terms of left cosets.
Proposition 2.
Let $X$ be an indecomposable left quasigroup, $x$ an element of $X$ and $H:=\mathcal{G}(X)_{x}$ the stabilizer of $x$. Then, $X$ is isomorphic to the left quasigroup on the left cosets $\mathcal{G}(X)/H$ given by $gH\cdotp hH:=\sigma_{g(x)}hH$ for all $g,h\in\mathcal{G}(X)$.
Proof.
By a standard calculation, one obtain that $(\mathcal{G}(X)/H,\cdotp)$ is a left quasigroup and $i:\mathcal{G}(X)/H\rightarrow X,$ $gH\mapsto g(x)$ is a left quasigroups isomorphism. ∎
The following result follows by a standard calculation.
Proposition 3.
Let $X$ be a latin left quasigroup. Then, $X$ is irretractable and $\mathrm{Dis}(X)$ is transitive. In particular, $X$ is indecomposable.
Proof.
Assume that $\sigma_{x}=\sigma_{y}$. Then $\delta_{z}(x)=x\cdotp z=y\cdotp z=\delta_{z}(y)$ for every $z\in X$. Thus $x=y$, i.e. $X$ is irretractable.
Since $X$ is latin, for every $a,t,z\in X$ there exist $b\in X$ such that $b\cdotp t=a\cdotp z$, therefore $\sigma_{b}^{-1}\sigma_{a}(z)=t$, hence $\mathrm{Dis}(X)$ is transitive on $X$. ∎
Given a left quasigroup $(X,\cdotp)$, we can define several operations that makes $X$ into a different left quasigroup.
Definition 2.
If $(X,\cdotp)$ is a left quasigroup and $\alpha\in Aut(X,\cdotp)$, we can define another operation $\cdotp_{\alpha}$ on $X$ by $x\cdotp_{\alpha}y:=\alpha(x\cdotp y)$ for every $x,y\in X$. The left quasigroup $(X,\cdotp_{\alpha})$ will be indicated by $X_{\alpha}$ and will be called deformation of $X$ by $\alpha$.
1.2 Coverings of indecomposable left quasigroups
In this section we develop a covering theory for indecomposable left quasigroup which is consistent with the one developed in [38, Section 2] for indecomposable cycle sets.
By Lemma 1, the following definition makes sense. Clearly, it is consistent with [38, Definition 1].
Definition 3.
Let $X,Y$ be indecomposable left quasigroups and $p$ an epimorphism from $X$ to $Y$. Then, $p$ is said to be a covering if the induced epimorphism $\bar{p}$ from $\mathcal{G}(X)$ to $\mathcal{G}(Y)$ is an isomorphism.
Note that if $p$ is a covering, then $\mathrm{Dis}_{\ker{p}}\leq\mathcal{G}^{\ker{p}}(X)=1$. Therefore,
$p(x)=p(y)\text{ implies }\sigma_{x}=\sigma_{y}.$ (2)
In the next result, we show that a left quasigroup $X$ always arises as epimorphic image of $\mathcal{G}(X)$ by a covering.
Proposition 4.
Let $X$ be an indecomposable left quasigroup, $x$ an element of $X$ and $Y$ be the left quasigroup $(\mathcal{G}(X),\cdotp)$ given by $g\cdotp h:=\sigma_{g(x)}h$ for all $g,h\in\mathcal{G}(X)$. Then the map
$u:Y\longrightarrow X,\quad g\mapsto g(x)$
is a epimorphism of left quasigroup. Moreover, $Y$ is a covering of $X$.
Proof.
It follows by a standard calculation using that $X$ can be identified as the left coset left quasigroup as in 2. ∎
Following the terminology of [38, Section 2] and [41, Section 3], $u$ will be called universal covering. Now, we give the analog of [41, Proposition 4] for arbitrary indecomposable left quasigroups.
Proposition 5.
Let $X,Y$ be indecomposable left quasigroups and suppose that $p:Y\rightarrow X$ is a covering. Then, the universal covering $u:\mathcal{G}(X)\rightarrow X$ factors through $p$.
Proof.
Let $x\in X$ and $y\in Y$ such that $p(y)=x$ and $\bar{p}:\mathcal{G}(Y)\longrightarrow\mathcal{G}(X)$ the induced group homomorphism, that is invertible. The map
$q:\mathcal{G}(X)\longrightarrow Y,\quad g\longrightarrow\bar{p}^{-1}(g)(y)$
is a left quasigroup epimorphisms and $u=p\circ q$. ∎
The following theorem, which is the main result of the section, is an extension of [38, Theorem 1] to indecomposable left quasigroups and allows to understand the structure of the epimorphisms.
Theorem 6.
Let $X,Y$ be indecomposable left quasigroups and $p$ an epimorphism from $X$ to $Y$. Then, $p$ factorize as $qr$ where $r:X\rightarrow X/\mathcal{G}^{\ker{p}}$ is the canonical projection and $q:X/\mathcal{G}^{\ker{p}}\rightarrow Y$, where $q$ is a covering of indecomposable left quasigroups.
Proof.
Let $\alpha=\mathcal{O}_{\mathcal{G}^{p}}$. Thus, the map $p$ factors through the canonical map onto $X/\alpha$,
[…]
Since $\alpha$ is contained in $\ker{p}$ then $\mathcal{G}^{\alpha}\leq\mathcal{G}^{\ker{p}}$. On the other hand $\mathcal{G}^{\ker{p}}\leq\mathcal{G}^{\alpha}$ since the blocks of $\alpha$ are the orbits of $\mathcal{G}^{\ker{p}}$. Thus, $\mathcal{G}^{\alpha}=\mathcal{G}^{\ker{p}}$. Therefore $\mathcal{G}(X/\alpha)=\mathcal{G}(X)/\mathcal{G}^{\alpha}=\mathcal{G}(X)/\mathcal{G}^{\ker{p}}=\mathcal{G}(Y)$. Therefore the map $q$ is a covering. ∎
According to (2), we obtain that in the irretractable case there are no non-trivial coverings.
Proposition 7.
Let $X,Y$ be left quasigroups and let $p:X\longrightarrow Y$ be a covering. If $X$ is irretractable then $p$ is an isomorphism.
Proof.
Note that $\ker{p}$ is contained in $\sim_{\sigma}=0_{Q}$. So $p$ is injective. ∎
1.3 Blocks of imprimitivity
Let $G$ be a group acting on a set $X$ and $\alpha$ be an equivalence relation of $X$. We say that $\alpha$ is $G$-invariant if $g(x)\,\alpha\,g(y)$ provided $x\,\alpha\,y$. If $G$ is transitive, a subset $Y\subseteq X$ is called a block of imprimitivity if $g(Y)\cap Y$ is either empty or equal to $Y$, for every $g\in G$, and the partition $P$ of $X$ given by $P:=\{g(Y)\}_{g\in G}$ is said to be a complete blocks system. If the action of $G$ has no non-trivial blocks of imprimitivity it is said to be primitive. If the action of $\mathcal{G}(X)$ is primitive we say that $X$ is primitive.
Let $X$ be a set and $H$ be a group of permutations on the set $X$. We define
$x\,\sim_{H}\,y\,\Longleftrightarrow\,H_{x}=H_{y}.$ (3)
$x\,\mathcal{O}_{H}\,y\,\Longleftrightarrow\,x=h(y),\text{ for some }h\in H.$ (4)
Clearly $\sim_{H}$ and $\mathcal{O}_{H}$ are equivalence relations.
Let us prove that the previous equivalence relations provide invariant partitions under further assumptions.
Lemma 8.
Let $G$ be a group acting on a set $X$ and $H\trianglelefteq G$. Then $\mathcal{O}_{H}$ and $\mathcal{O}_{H}$ and $\sim_{H}$ are $G$-invariant equivalence relation on $X$.
Proof.
Let $g\in G$, $h\in H$ and $x\in X$. Then $gh(x)=ghg^{-1}g(x)$. Then $g(x)$ and $gh(x)$ are in the same orbit w.r.t. $G$.
If $H_{x}=H_{y}$ then $H_{g(x)}=gH_{x}g^{-1}=gH_{y}g^{-1}=H_{g(y)}$, since $H$ is normal in $G$. ∎
Corollary 9.
Let $X$ be a left quasigroup and $H$ be a normal subgroup of $\mathcal{G}(X)$. Then $\sim_{H}$ is a $\mathcal{G}(X)$-invariant equivalence relation.
Proof.
We can apply Lemma 8. ∎
Note that Corollary 9 applies to admissible subgroups.
Lemma 10.
Let $X$ be a left quasigroup and $H$ be a normal subgroup of $\mathcal{G}(X)$.
  • (i)
    If $\sim_{H}=0_{X}$ then $N_{\mathcal{G}(X)}(H_{x})\leq\mathcal{G}(X)_{x}$ for every $x\in X$. In particular $Z(\mathcal{G}(X))=1$.
  • (ii)
    $\sim_{H}=1_{X}$ if and only if $H_{x}=1$ for every $x\in X$, i.e. $H$ is semiregular.
Proof.
Let $G:=\mathcal{G}(X)$.
(i) Let $g\in N_{G}(G_{x})$. Then $H_{x}=gH_{x}g^{-1}=H_{g(x)}$. Then $x\,\sim_{H}\,g(x)$ and so $g(x)=x$. Namely, $N_{G}(G_{x})\leq G_{x}$. Since $Z(G)\leq C_{G}(H_{x})$ for every $x\in X$, then $Z(G)\leq\cap_{x\in X}H_{x}=1$.
(ii) Let $g\in H_{x}$. If $x\,\sim_{H}\,y$ for every $x,y\in X$ then $g(y)=y$ for every $y\in X$, i.e. $g=1$. ∎
Corollary 11.
Let $X$ be a indecomposable left quasigroup and $H$ be a normal subgroup of $\mathcal{G}(X)$. Then $\sim_{H}=0_{X}$ if and only if $N_{\mathcal{G}(X)}(H_{x})\leq\mathcal{G}(X)_{x}$ for every $x\in X$.
Proof.
The stabilizers are all conjugate since for every $x,y\in X$ there exists $g\in\mathcal{G}(X)$ such that $y=g(x)$. Therefore $H_{y}=H_{g(x)}=gH_{x}g^{-1}=H_{x}$ if and only if $g\in N_{\mathcal{G}(X)}(H_{x})$. Therefore, $\sim_{H}=1$ if and only if $N_{\mathcal{G}(X)}(H_{x})\leq\mathcal{G}(X)_{x}$ for every $x\in X$. ∎
Note that congruences of a left quasigroup $X$ are $\mathcal{G}(X)$-invariant. We show that the converse does not hold in general by using the equivalence relations defined in this section.
Example 12.
Let $X:=\{1,2,3,4\}$ be the left quasigroup given by $\sigma_{1}:=(2,4)$, $\sigma_{2}:=(1,2,3,4)$, $\sigma_{3}:=(1,4,3,2)$, and $\sigma_{4}:=(1,3)$. Then, $\mathcal{G}(X)$ is isomorphic to the dihedral group of size $8$ and the subgroup $H$ generated by the set $\{(1,3)(2,4),(2,4)\}$ is a normal subgroup of $\mathcal{G}(X)$. The relations $\sim_{H}$ and $\mathcal{O}_{H}$ coincide and are given by the partition $\{\{1,3\},\{2,4\}\}$. Such relations are not congruences of $X$.
2 Left braces
Following [15, Definition 1], a set $B$ endowed of two operations $+$ and $\circ$ is said to be a left brace if $(B,+)$ is an abelian group, $(B,\circ)$ a group, and
$x\circ(y+z)+x=x\circ y+x\circ z,$
for all $x,y,z\in B$. The group $(B,+)$ will be called the additive group of the left brace, while the group $(B,\circ)$ will be called the multiplicative group.
Examples 13.
  • 1)
    If $(B,+)$ is an abelian group, then the operation $\circ$ given by $x\circ y:=x+y$ give rise to a left brace which we will call trivial.
  • 2)
[…]
If $(B_{1},+,\circ)$ and $(B_{2},+^{\prime},\circ^{\prime})$ are left braces, a homomorphism $\psi$ between $B_{1}$ and $B_{2}$ is a function from $B_{1}$ to $B_{2}$ such that $\psi(x+y)=\psi(x)+^{\prime}\psi(y)$ and $\psi(x\circ y)=\psi(x)\circ^{\prime}\psi(y)$, for all $x,y\in B_{1}$.
Given a left brace $B$ and $x\in B$, let us denote by $\lambda_{x}:B\longrightarrow B$ the map from $B$ into itself defined by
$\lambda_{x}(y):=-x+x\circ y,$ (5)
for all $y\in B$. Let us recall the properties of the maps $\lambda_{x}$.
Proposition 14.
[35, Proposition 2],[15, Lemma 1] Let $B$ be a left brace. Then, the following are satisfied:
  • 1)
    $\lambda_{x}\in\operatorname{Aut}(B,+)$, for every $x\in B$;
  • 2)
    the map $\lambda:B\longrightarrow\operatorname{Aut}(B,+)$, $x\mapsto\lambda_{x}$ is a group homomorphism from $(B,\circ)$ into $\operatorname{Aut}(B,+)$.
For the following definition, we refer the reader to [35, pg. 160] and [15, Definition 3].
Definition 4.
Let $B$ be a left brace. A subset $I$ of $B$ is said to be a left ideal if it is a subgroup of the multiplicative group and $\lambda_{x}(I)\subseteq I$, for every $x\in B$. Moreover, a left ideal is an ideal if it is a normal subgroup of the multiplicative group.
If $I$ is an ideal of a left brace $B$, the quotient set $B/I$ has a canonical left brace structure called the quotient left brace of $B$ modulo $I$ (the operations are defined as $(x+I)+(y+I)=x+y+I$ and $(x+I)\circ(y+I)=(x\circ y)+I)$ for every $x,y\in B$). Ideals are nothing other than kernels of homomorphisms, where the kernel of a homomorphism $\psi$ is the set given by $\ker(\psi):=\{x\in B_{1}|\psi(x)=0\}$. Indeed, the kernel of a homomorphism is always an ideal of a left brace, and conversely an ideal $I$ is the kernel of the canonical projection $B\longrightarrow B/I$.
The set $\{0\}$ is an ideal and it will be called the trivial ideal. A left brace $B$ which contains no ideals different from $\{0\}$ and $B$ will be called a simple left brace.
A standard ideal of a left brace $B$ is the socle, indicated by $Soc(B)$ and given by the kernel of the map $\lambda$. Another useful ideal of a left brace $B$, given in [35, Corollary of Proposition 6] and indicated by $B^{2}$, is the one given by the additive subgroup generated by the set $\{x*y\hskip 2.84526pt|\hskip 2.84526ptx,y\in B\}$, where $x*y:=-x+x\circ y-y$ for all $x,y\in B$. In general, if $I$ is an ideal of a left brace $B$, then $I*B$ and $B*I$ are subset of $I$, where $I*B$ and $B*I$ are defined in the same way of $B^{2}$.
In general, if $B$ is a left brace and $x\in B$ the map $\lambda_{x}$ does not belong to $Aut(B,+,\circ)$. The following lemma, that contains the same idea of [22, Lemma $1.1$], shows that under additional hypothesis, a map $\lambda_{x}$ induces an automorphism on some left ideals.
Lemma 15.
Let $B$ be a left brace and $C,D$ two left ideals of $B$. Moreover, suppose that $\lambda_{x}(y)=y$ for all $x\in C+D$ and $y\in D$. Then, the restriction of $\lambda_{y}$ to $C$ is an element of $Aut(C,+,\circ)$.
Proof.
Let $y\in D$ and $z\in C.$ Then we have
$\lambda_{y}(z)=-y+y\circ z=-y+y\circ z\circ y^{-}\circ y=$
$=$ $-y+y\circ z\circ y^{-}-y\circ z\circ y^{-}+y\circ z\circ y^{-}\circ y$
$=$ $-y+y\circ z\circ y^{-}+\lambda_{y\circ z\circ y^{-}}(y)$
$=$ $y\circ z\circ y^{-}$
therefore the restriction of $\lambda_{y}$ to $C$ is an element of $Aut(C,+,\circ)$. ∎
Recall that every Sylow subgroup of the additive group is always a left ideal. Using the left ideal structures, one can show that if $B_{1}$,…,$B_{t}$ are Sylow subgroups of $(B,+)$, then the subgroup $B_{1}+...+B_{t}$ of the additive group is equal to $B_{1}\circ...\circ B_{t}$ and is also a subgroup of the multiplicative group.
Lemma 16.
(Proposition $3$, [37] and Proposition $2.6$, [16]) Let $B$ be a left brace and $B_{p_{1}},...,B_{p_{r}}$ be the Sylow subgroups of $(B,+)$. Then:
  • 1)
    if $N$ is a normal subgroup of $(B,\circ)$ contained in $B_{p_{1}}$, then $\lambda_{n}(b)=b$ for all $n\in N$ and $b\in B_{p_{2}}\circ...\circ B_{p_{r}}$;
  • 2)
    $B_{p_{1}}$ is a normal subgroup of $(B,\circ)$ if and only if $\lambda_{g}(b)=b$ for all $g\in B_{p_{1}}$ and $b\in B_{p_{2}}\circ...\circ B_{p_{r}}$.
3 Cycle sets
3.1 Basics
Following [34], a left quasigroup $(X,\cdot)$ is said to be a cycle set if the equality
$(x\cdot y)\cdot(x\cdot z)=(y\cdot x)\cdot(y\cdot z),$ (6)
holds for all $x,y,z\in X$. Moreover, a cycle set $(X,\cdot)$ is called non-degenerate if the squaring map $\mathfrak{q}:X\longrightarrow X$, $x\mapsto x\cdot x$ is bijective. A cycle set will be called irretractable/indecomposable/latin if its underlying left quasigroup is irretractable/indecomposable/latin.
Examples 17.
  • 1)
    If $X$ is a nonempty set and $\gamma\in Sym(X)$, the binary operation given by $x\cdotp y:=\gamma(y)$ makes $X$ into a non-degenerate cycle set. We will refer to these cycle sets as the trivial cycle set provided by the permutation $\gamma$.
  • 2)
    If $X$ is a cycle set, then every deformation of $X$ is again a cycle set (see [40, Section $3$]).
Remarks 1.
  • 1)
    From now on, every cycle set will be non-degenerate. All the results will be given in the language of cycle sets, but they can be translated in terms of solutions by Eq. 1.
  • 2)
    A notion of displacement group for cycle sets was given in [9, Section $2$]. Even if it is slightly different from the one given in Section $1$ for arbitrary left quasigroup, by [9, Lemma $2.15$ and Proposition $2.17$] these two definitions coincide for non-degenerate cycle sets.
Let $B$ be a left brace. A cycle base is a subset $X$ of $B$ that is a union of orbits with respect to the $\lambda$-action and generating the additive group $(B,+)$. If a cycle base is a single orbit, then is called transitive cycle base. By [38, Theorem 3], every indecomposable cycle sets can be constructed by transitive cycle bases of left braces; on the other hand, every cycle set $X$ gives rise to a left brace structure on $\mathcal{G}(X)$ having as $\circ$ operation the usual maps composition (see Section $1$ of [38]).
Proposition 18 (Theorem 3, [38]).
Let $(B,+,\circ)$ be a left brace, $Y\subset B$ a transitive cycle base, $a\in Y$, and $K$ a core-free subgroup of $(B,\circ)$, contained in the stabilizer $B_{a}$ of $a$ (respect to the action $\lambda$). Then, the pair $(X,\cdotp)$ given by $X:=B/K$ and $\sigma_{x\circ K}(y\circ K):=\lambda_{x}(a)^{-}\circ y\circ K$ give rise to an indecomposable cycle set with $\mathcal{G}(X)\cong B$.
Conversely, every indecomposable cycle set $(X,\cdotp)$ can be obtained in this way, taking $B:=\mathcal{G}(X)$.
The following result, implicitly contained in [38], allows to avoid the construction of the previous proposition to recover indecomposable and irretractable cycle sets.
Proposition 19 (proof of Theorem 3, [38]).
Let $X$ be an indecomposable and irretractable cycle set. Then, there exist a transitive cycle base $Z$ of $\mathcal{G}(X)$ such that $X$ is isomorphic, as cycle set, to the cycle set on $Z$ given by $x\cdotp y:=\lambda_{x}^{-1}(y)$ for all $x,y\in Z$.
By [41, Section 3], we know that if $X$ is an indecomposable cycle set and $I$ is an ideal of the left brace $\mathcal{G}(X)$, the $I$-orbits on $X$ give rise to a congruence of the cycle set $X$. This result can be recovered by looking at the connection between ideals and admissible subgroups.
Proposition 20.
Let $X$ be an indecomposable cycle set and $H$ be an ideal of $\mathcal{G}(X)$. Then $H$ is an admissible subgroup. If $X$ is finite, the converse also follows.
Proof.
Let $H$ be an ideal of the left brace $\mathcal{G}(X)$. Then $H$ is a normal subgroup of the multiplicative group. Moreover,
$\sigma_{x}\circ\sigma^{-1}_{h(x)}=\sigma_{x}\circ\lambda_{h}(\sigma_{x}^{-1})=\sigma_{x}\circ h\circ(h^{-1}+\sigma_{x}^{-1})$
$=$ $\sigma_{x}\circ h\circ\sigma_{x}^{-1}\circ\sigma_{x}\circ(h^{-1}+\sigma_{x}^{-1})$
$=$ $\sigma_{x}\circ h\circ\sigma_{x}^{-1}\circ\lambda_{\sigma_{x}}(h)$
for all $x\in X,h\in H$. Since $\sigma_{x}\circ h\circ\sigma_{x}^{-1}\in H$ and $\lambda_{\sigma_{x}}(h)\in H$ for all $x\in X,h\in H$, we have that $H$ is an admissible subgroup of $\mathcal{G}(X)$.
If $X$ is finite and $H$ is an admissible subgroup, as in the previous implication we obtain $\sigma_{x}\circ\sigma^{-1}_{h(x)}=\sigma_{x}\circ h\circ\sigma_{x}^{-1}\circ\lambda_{\sigma_{x}}(h)$ for all $x\in X$, $h\in H$. Since $H$ is a normal subgroup, we have that $\lambda_{\sigma_{x}}(h)$ for all $x\in X,h\in H$, and by finiteness of $X$, since the $\sigma_{x}$ generates $\mathcal{G}(X)$, we have that $\lambda_{g}(h)\in H$ for all $g\in\mathcal{G}(X),h\in H$. Hence $H$ is an ideal of $\mathcal{G}(X)$. ∎
Corollary 21 (Section 3, [41]).
Let $X$ be a cycle set and $I$ an ideal of $\mathcal{G}(X)$. Then $\mathcal{O}_{I}$ is a congruence of $X$.
Proof.
The statement follows by 20 and [7, Lemma 1.8]. ∎
3.2 Indecomposable cycle sets of prime-power size
In this section we consider irretractable and indecomposable cycle sets having prime-power order. Under an additional hypothesis, these cycle sets can be obtained by irretractable and indecomposable cycle sets with a $p$-group permutation group.
Theorem 22.
Let $X$ be an indecomposable and irretractable cycle set of size $p^{n}$ for some prime number $p$. Let $B_{p}$ the $p$-Sylow subgroup of $(\mathcal{G}(X),+)$. The following are equivalent:
  1. 1.
    $B_{p}$ is a normal subgroup of $(\mathcal{G}(X),\circ)$.
  2. 2.
    $X$ is a deformation $Y_{\alpha}$ of a cycle set $Y$ such that $\mathcal{G}(Y)\cong B_{p}$, $\alpha$ has order coprime with $p$ and $\alpha$ fixes an element $y\in Y$.
Proof.
Assume that $B_{p_{1}}:=B_{p}$ is a normal subgroup of $(\mathcal{G}(X),\circ)$ and let $B_{p_{2}},...,B_{p_{r}}$ be the Sylow subgroup of $(\mathcal{G}(X),\circ)$ different from $B_{p}$. If $B_{p}=\mathcal{G}(X)$, we can take $\alpha=id_{X}$ to obtain the statement. Let $B_{p}<\mathcal{G}(X)$. Since $X$ is irretractable, we can identify $X$ with a transitive cycle base of $\mathcal{G}(X)$. By orbit-stabilizer equation, if $St(x)$ is the stabilizer of $x$ respect to the action $\lambda$, we have $|\mathcal{G}(X)|=p^{n}\cdotp|St(x)|$. Since $\mathcal{G}(X)$ is solvable and finite, we have that the subgroup $B_{p_{2}}\circ...\circ B_{p_{r}}$ is conjugate to a subgroup of $St(x)$ so, up to replace $x$, we can suppose the inclusion $B_{p_{2}}\circ...\circ B_{p_{r}}\subseteq St(x)$. Now, let $x_{1},...,x_{r}\in\mathcal{G}(X)$, with $x_{i}\in B_{p_{i}}$ for every $i\in\{1,...,r\}$, such that $x=x_{1}+...+x_{r}$. Since this decomposition is unique and the subgroups $B_{p_{i}}$ are left ideals, it follows that $\lambda_{a}(x_{i})=x_{i}$ for all $a\in B_{p_{2}}\circ...\circ B_{p_{r}}$ and $i\in\{1,...,r\}$. Moreover, being $B_{p}$ normal in $\mathcal{G}(X)$, by Lemma 16 we have $\lambda_{e}(x_{i})=x_{i}$ for all $e\in B_{p}$ and $i\in\{2,...,r\}$. From these facts, it follows that $\lambda_{g}(x_{i})=x_{i}$ for all $g\in\mathcal{G}(X)$ and $i\in\{2,...,r\}$. Since $X$ is identified with a $\lambda$-orbit, if $y,z$ are elements of $X$, there exist $y_{1},z_{1}\in B_{p_{1}}$ such that $y=y_{1}+t$ and $z=z_{1}+t$, where $t=x_{2}+...+x_{r}$. Then
$y\cdotp z=\lambda_{y_{1}+t}^{-1}(z_{1}+t)=\lambda^{-1}_{\lambda^{-1}_{y_{1}}(t)}(\lambda^{-1}_{y_{1}}(z_{1}))+\lambda^{-1}_{\lambda^{-1}_{y_{1}}(t)}(\lambda^{-1}_{y_{1}}(t))$
and since $\lambda_{g}(t)=t$ for all $g\in\mathcal{G}(X)$, we obtain
$y\cdotp z=\lambda_{t}^{-1}(\lambda^{-1}_{y_{1}}(z_{1}))+t.$ (7)
Therefore, we have that $X$ is the cycle set given by the lambda orbit $Y$ of $x_{1}$ respect to $B_{p}$ composed with $\lambda_{t}^{-1}$. By Lemma 15, we obtain easily that the restriction of $\lambda_{t}^{-1}$ to $Y$ is a cycle set automorphism, hence $X$ is a deformation of $Y$ by $\lambda_{t}^{-1}$, which has order coprime with $p$ and fixes $x_{1}$, and $\mathcal{G}(Y)\cong B_{p}$.
The converse follows by [22, Lemma 4.5]. ∎
Remark 1.
The previous result reduces the description of all indecomposable and irretractable cycle sets of prime-power size $p^{n}$ and with normal Sylow $p$-subgroup permutation group to the ones having a $p$-group permutation group. This is consistent with the first part of [22, Section $4$]. However, we highlight that 22 is not useful to describe all the irretractable and indecomposable cycle set of size $p^{n}$. Indeed, Example $5$ of [41] is an irretractable and indecomposable cycle set of size $8$ and with permutation group isomorphic to $\mathbb{S}_{4}$.
Now, we can determine the structure of the permutation left braces associated to the cycle sets of 22. Recall that in a left brace $B$ the Sylow subgroups of $(B,+)$ coincide with the ones of $(B,\circ)$, and moreover if $B_{1},...,B_{t}$ are Sylow subgroup of $(B,+)$, then $B_{1}+...+B_{t}=B_{1}\circ...\circ B_{t}$. Now, if $X$ is an indecomposable and irretractable cycle set of prime-power size $p^{n}$, we indicate by $B_{p_{1}}:=B_{p}$, $B_{p_{2}},...,B_{p_{r}}$ the Sylow subgroups of $\mathcal{G}(X)$.
Proposition 23.
If $B_{p_{1}}$ is normal in $\mathcal{G}(X)$, then $\mathcal{G}(X)$ is isomorphic, as left brace, to the left braces semidirect product $B_{p_{1}}\rtimes_{\alpha}(B_{p_{2}}+...+B_{p_{r}})$, where $B_{p_{2}}+...+B_{p_{r}}$ is a trivial left brace with cyclic additive group and $\alpha$ correspond to the action by the map $\lambda$.
Proof.
It follows by 22 and [22, Lemma 4.5]. ∎
As a corollary, we recover [25, Theorem 2.13], that classify indecomposable cycle sets of prime size.
Corollary 24.
(Theorem 2.13, [25]) Let $X$ be an indecomposable cycle set of prime size $p$. Then $X$ is isomorphic to the deformation of the trivial cycle set of size $p$ by an arbitrary $p$-cycle.
Proof.
Suppose that $X$ is irretractable. Since by [25, Theorem $2.15$] $\mathcal{G}(X)$ is solvable, by [33, Exercise $7.2.12$] we have that it is isomorphic to a subgroup of the form $\mathbb{Z}/p\mathbb{Z}\rtimes N$, with $N$ a cyclic subgroup having order dividing $p-1$. Therefore, by 22 $X$ is a deformation of an indecomposable cycle set $\bar{X}$ with permutation group isomorphic to $\mathbb{Z}/p\mathbb{Z}$. Since $X$ is irretractable, it follows that $\bar{X}$ is irretractable, but this contradicts [14, Proposition 1]. Hence, we have that $X$ is retractable. Therefore by [12, Lemma 1] $X$ is a cycle set given by $x\cdotp y:=\alpha(y)$ for all $x,y\in X$, where $\alpha$ is a $p$-cycle. By a standard calculation, one can show that the isomorphism class of this cycle set does not depend on the choice of the $p$-cycle. ∎
3.3 Blocks of imprimitivity
We turn our attention on the primitive cycle sets, i.e. cycle sets having primitive permutation group. In this subsection, we recover the classification of these cycle sets given in [16], and we provide a forward improvement that give us a method to find a non-trivial complete blocks system.
At first, recall that a permutation group $G$ of a finite set $X$ is said to be quasiprimitive if every normal subgroup acts transitively on $X$ and it is called innately transitive if there exist a minimal normal subgroup that acts transitively on $X$. Clearly, the implications “primitive $\Rightarrow$ quasiprimitive” and “quasiprimitive $\Rightarrow$ innately transitive” hold. For solvable groups, these notions are equivalent (for more details, see [6, Section 1] and [31]).
Lemma 25.
Let $G$ be a solvable permutation group of a finite set $X$. Then, the following statements are equivalent:
  • 1)
    $G$ acts primitively on $X$;
  • 2)
    $G$ acts quasiprimitively on $X$;
  • 3)
    the action of $G$ on $X$ is innately transitive.
Proposition 26.
Let $X$ be an indecomposable cycle set with innately transitive permutation group. Then, $X$ is a trivial indecomposable cycle set of size $p$, for some prime number $p$.
Proof.
By [25, Theorem 2.15], $\mathcal{G}(X)$ is a finite solvable group, hence by Lemma 25 we have that $\mathcal{G}(X)$ acts primitively on $X$. Again by solvability of $\mathcal{G}(X)$ and by [33, Theorem 7.2.6] we have that $|X|=p^{n}$ for some natural number $n$ and $\mathcal{G}(X)$ is isomorphic to the semidirect product $N\rtimes G$ where $N$ is an elementary abelian $p$-group that acts transitively on $X$ and $G$ is the stabilizer of an element $x\in X$. Moreover, $N$ is a minimal normal subgroup of $\mathcal{G}(X)$ and is self-centralizing. By primitivity of $\mathcal{G}(X)$, either $X$ is irretractable or $n$ is equal to $1$. Suppose that $X$ is irretractable. Since $\mathcal{G}(X)$ is solvable and finite, we can suppose that $G$ contains $B_{p_{2}}+...+B_{p_{r}}$ and by 19, we have that $X$ can be identified with a transitive cycle base, therefore $\lambda_{a}(x)=x$ for all $a\in G$ and, if $x_{1}\in B_{p_{1}}$,…,$x_{r}\in B_{p_{r}}$ are such that $x=x_{1}+...+x_{r}$, it follows that $\lambda_{a}(x_{i})=x_{i}$ for all $i\in\{1,...,r\}$. Moreover, since $N$ is a normal $p$-group, by Lemma 16 it follows that $\lambda_{n}(x_{i})=x_{i}$ for all $i\in\{2,...r\}$. By these facts, we obtain that $B_{p_{2}}+...+B_{p_{r}}$ is a trivial left brace with cyclic additive group, and that $B_{p}$ acts trivially on $B_{p_{2}}+...+B_{p_{r}}$ by the map $\lambda$, therefore by Lemma 16 $B_{p}$ is normal in $\mathcal{G}(X)$. By minimality and self-centrality of $N$, it follows that $N=Z(B_{p})=B_{p}$. By 22, $X$ is a deformation of an irretractable cycle set $\bar{X}$ with abelian permutation group isomorphic to $N$, but this contradicts [14, Proposition 1]. Therefore, $X$ has prime size $p$. The rest of the statement follows by 24. ∎
As we said before, 26 is essentially the main result of [16] with a forward improvement. The main difference is that now we have a concrete method to find at least a complete blocks system by the relation given in Eq. 4.
Corollary 27.
Let $X$ be a cycle set. Then, $X$ is primitive if and only if it is isomorphic to a cycle set of prime size $p$ given by $(\mathbb{Z}/p\mathbb{Z},\cdotp)$ and $x\cdotp y:=y+1$ for all $x,y\in\mathbb{Z}/p\mathbb{Z}$
4 Cycle sets with regular displacement group
4.1 Right linear cycle sets
Let us introduce a family of left quasigroups constructed over groups.
Definition 5.
A left quasigroup $X$ is said to be right linear over a group $(G,\circ)$ if $X=G$ and there exist a map $\phi:G\longrightarrow G$ and $\psi\in Aut(G,\circ)$ such that $x\cdotp y:=\phi(x)\circ\psi(y)$ for all $x,y\in X$. If in addition $\phi\in Sym_{G}$, $X$ is latin.
If $(G,\circ)$ is abelian and $x\cdot y=\phi(x)+\psi(y)+c$ for $\phi\in\operatorname{End}(G,+)$ and $c\in G$ we say that $(G,\cdotp)$ is affine. We denote such a cycle set by $\mathrm{Aff}(G,\phi,\psi,c)$.
By a standard calculation, one can show that a left quasigroup $X$ right linear over a group $(G,+)$ is a cycle set if and only if
$\phi(\phi(y)+\psi(x))+\psi(\phi(y))=\phi(\phi(x)+\psi(y))+\psi(\phi(x))$ (8)
holds for all $x,y\in G$. In a similar way, it follows that a left quasigroup $X$ affine over a group $(G,+)$, where $x\cdotp y=\phi(x)+\psi(y)+c$ for a constant $c\in G$, is a cycle set if and only if $\phi\psi(x)-\psi\phi(x)=\phi^{2}(x)$ for all $x\in G$. The automorphism group of affine latin left quasigroups are completely described.
Lemma 28.
([23, Proposition $3.9$]) Let $X=\mathrm{Aff}(G,\phi,\psi,c)$ be a latin left quasigroup affine over an abelian group $(G,+)$. Then
$Aut(G,\cdotp)=\{t_{g}\circ\eta\mid\eta\in C_{Aut(G,+)}(\phi,\psi),\quad\eta(c)-c=\phi(g)+\psi(g)-g\}$
Deformations preserve in several cases the property of being affine or right linear, as we can see in the following proposition.
Proposition 29.
Let $X$ be a left quasigroup, $\alpha\in\operatorname{Aut}{(X,\cdot)}$ and $G$ be a group. Then:
  • 1)
    $X$ is irretractable (resp. latin) if and only if $X_{\alpha}$ is irretractable (resp. latin).
  • 2)
    $\mathrm{Dis}(X)=\mathrm{Dis}(X_{\alpha})$.
  • 3)
    If $G$ is abelian then $X$ is affine over $G$ if and only if $X_{\alpha}$ is affine over $G$.
  • 4)
    If $X$ is a latin cycle set, then $X$ is right linear over $G$ if and only if $X_{\alpha}$ is right linear over $G$.
Proof.
1) Let $x,y\in X$ and let $\bar{\sigma}_{x}$, $\bar{\delta}_{x}$ denote the left and right multiplication by $x$ in $X_{\alpha}$. Note that $\bar{\sigma}_{x}=\alpha\sigma_{x}$ and $\bar{\delta}_{x}=\alpha\delta_{x}$.
Clearly $\sigma_{x}=\sigma_{y}$ if and only if $\bar{\sigma}_{x}=\bar{\sigma}_{y}$ and $\delta_{x}$ is bijective if and only if $\bar{\delta}_{x}$ is bijective. Hence $X$ is irretractable (resp. latin) if and only if $X_{\alpha}$ is irretractable (resp. latin).
2) Since $\alpha\in Aut(X)$, we obtain $\bar{\sigma}_{x}\bar{\sigma}^{-1}_{y}=\alpha\sigma_{x}\sigma_{y}^{-1}\alpha^{-1}=\sigma_{\alpha(x)}\sigma^{-1}_{\alpha(y)}$, hence $\mathrm{Dis}(X)=\mathrm{Dis}(X_{\alpha})$.
3) If $X$ is a quasigroup affine over an abelian group $G$ and $\alpha\in Aut(X,\cdotp)$, by Lemma 28 there exist $\eta\in Aut(G,\circ)$ and $g\in G$ such that $x\cdotp_{\alpha}y=g\circ\eta(\phi(x)\circ\psi(y)\circ c)$ for all $x,y\in X$, and therefore $x\cdotp_{\alpha}y=\eta\phi(x)\circ\eta\psi(y)\circ(g\circ\eta(c))$ for all $x,y\in X$. Hence $X_{\alpha}$ is affine over $(G,\circ)$. Since $\alpha^{-1}\in Aut(X,\cdotp_{\alpha})$ and $(X,\cdotp)$ can be obtained by the deformation of $X_{\alpha}$ by $\alpha^{-1}$, the converse also follows.
4) By 1), 2) and [9, Theorem $5.3$], we have that if $X$ is a latin cycle set right linear over a group $G$, then so is $X_{\alpha}$. The converse follows as in 3). ∎
Definition 6.
Let $G$ be a group, $H$ a normal subgroup of $G$ and $f$ a map from $G$ to itself. Then, $H$ is said to be invariant under $f$ if $x^{-1}y\in H$ implies that $f(x)^{-1}f(y)\in H$ for every $x,y\in G$.
Let us show a description of imprimitive blocks systems and congruences of right linear left quasigroups.
Proposition 30.
Let $X$ be a finite latin left quasigroup right linear over a group $(G,\circ)$. Then, the complete blocks systems of $\mathcal{G}(X)$ are in one-to-one correspondence with the subgroups of $G$ that are invariant under $\psi$.
Proof.
Let $H$ be a subgroup of $G$ invariant under $\psi$ and set $B_{x,H}:=\{x\circ h\mid h\in H\}$ for all $x\in G$. Then, by a standard calculation one can show that the set $\mathcal{B}_{H}:=\{B_{x,H}\}_{x\in G}$ is a complete blocks system.
Conversely, let $\mathcal{B}$ be a complete blocks system and let $B$ be the element of $\mathcal{B}$ such that $0\in B$. Let $y=\phi^{-1}(0)$. Then $\sigma_{y}=\psi$ and so $\sigma_{y}(0)=0$ and accordingly $\psi(B)=\sigma_{y}(B)=B$.
We indicate by $\sim$ the equivalence relation induced by $\mathcal{B}$. At first, we show that $B$ is a subgroup of $G$. Now, let $y_{1},y_{2}\in B$. Then, since $\phi$ is a bijection, we have that $s\circ\psi(y_{1})\sim s\circ\psi(y_{2})$ for all $s\in G$. If we set $s:=\psi(y_{1})^{-}$, since $\psi\in Aut(G,\circ)$ we obtain $\psi(y_{1}^{-}\circ y_{2})\sim 0$, therefore $\psi(y_{1}^{-}\circ y_{2})\in B$, and hence $y_{1}^{-}\circ y_{2}\in B$. Therefore, it follows that $B$ is a subgroup of $G$. By a standard calculation, we obtain that $\mathcal{B}=\mathcal{B}_{B}$, hence the statement follows. ∎
The following result improves [9, Lemma 4.8] in the finite case.
Proposition 31.
Let $X$ be a finite latin left quasigroup right linear over a group $(G,\circ)$. Then, the congruence of $X$ are in one-to-one correspondence with normal subgroups $H$ of $G$ that are invariant under $\psi$ and $\phi$.
Proof.
Let $\sim$ be a congruence of $X$. By 30, there exist a $\psi$-invariant subgroup $H$ of $G$ such that $\sim$ coincides with the equivalence relation $\sim_{H}$ induced by $H$. Now, if $x\sim_{H}x^{\prime}$ we have that $x\cdotp 0\sim x^{\prime}\cdotp 0$, hence $\phi(x)\sim_{H}\phi(x^{\prime})$ and moreover $\phi^{-1}(0)\cdotp x\sim_{H}\phi^{-1}(0)\cdotp x$, and therefore $H$ is invariant under $\phi$. From the finitneness of $X$ it follows that $\phi^{k}(x)\sim_{H}\phi^{k}(x^{\prime})$ and $\psi^{k}(x)\sim_{H}\psi^{k}(x^{\prime})$ for all $k\in\mathbb{Z}$. Now, let $x,x^{\prime},y,y^{\prime}\in X$ such that $x\sim_{H}x^{\prime}$ and $y\sim_{H}y^{\prime}$. Then, since $\sim_{H}$ is a congruence of $X$, we have $\phi^{-1}(x)\cdotp\psi^{-1}(y)\sim_{H}\phi^{-1}(x^{\prime})\cdotp\psi^{-1}(y^{\prime})$ and hence $x\circ y\sim_{H}x^{\prime}\circ y^{\prime}$, i.e. $\sim_{H}$ is a congruence of $(G,\circ)$. Therefore $H$ is a normal subgroup.
The converse is a standard calculation. ∎
4.2 Right linear representation
A left quasigroup $X$ has semiregular displacement group if and only if $\sim_{\mathrm{Dis}(X)}=1_{X}$ (see Eq. 3). Left quasigroups with semiregular displacement groups have been studied in [8], and in particular we have the following corollary of [8, Lemma 3.2].
Proposition 32.
Every finite irretractable left quasigroup with semiregular displacement group is a latin.
The connection between right linear left quasigrups and cycle sets has been studied in [9], in particular, we have the following reformulation of [9, Theorem $5.3$], since finite irretractable cycle sets with semiregular displacement group are latin.
Corollary 33.
For a finite irretractable cycle set $X$ the following conditions are equivalent:
  • 1)
    $X$ is right linear over a group $G$ with $G\cong\mathrm{Dis}(X)$.
  • 2)
    $\mathrm{Dis}(X)$ acts regularly on $X$.
Proof.
If $X$ is an irretractable cycle set, by a standard calculation we have that if either 1)1)1 ) or 2)2)2 ) hold, then $X$ is latin. Hence, the statement follows by [9, Theorem 5.3]. ∎
Remark 2.
If $(X,\cdotp)$ is an irretractable cycle set with regular displacement group, by 33 we can represent it as a cycle set right linear over a suitable group $(G,\circ)$ and suitable maps $\phi$ and $\psi$ as in 5. The theory contained in [9, Section 5] gives a method to find the operation $\circ$ and the maps $\phi$ and $\psi$, even if these maps are not uniquely determined. Indeed, if we choose two elements $e,f\in X$, then $X$ has a group structure given by $x\circ_{e,f}y:=\delta_{e}^{-1}(x)\cdotp\sigma_{f}^{-1}(y)$ for all $x,y\in X$ , where $f\cdotp e$ is the neutral element. In this way, $\phi_{e,f}(x):=\delta_{e}(x)\circ_{e,f}\sigma_{f}(e\cdotp f)$ and $\psi_{e,f}(x):=\sigma_{f}(e\cdotp f)^{-1}\circ_{e,f}\sigma_{f}(x)$ for all $x\in X$. The displacement group $\mathrm{Dis}(X)$ is isomorphic to $(X,\circ_{e,f})$, and an explicit isomorphism is given by $\rho_{e,f}:(X,\circ_{e,f})\rightarrow\mathrm{Dis}(X)$, $x\mapsto\sigma_{\delta_{e}^{-1}(x)}\sigma_{f}^{-1}$ for all $x\in X$. In this way, the map $\psi_{e,f}$ induces an automorphism of $\mathrm{Dis}(X)$ given by $\overline{\psi_{e,f}}(\sigma_{\delta_{e}^{-1}(x)}\sigma_{f}^{-1}):=\sigma_{\delta_{e}^{-1}(\psi_{e,f}(x))}\sigma_{f}^{-1}$ and the map $\phi_{e,f}$ induces a bijection given by $\overline{\phi_{e,f}}(\sigma_{\delta_{e}^{-1}(x)}\sigma_{f}^{-1}):=\sigma_{\delta_{e}^{-1}(\phi_{e,f}(x))}\sigma_{f}^{-1}$, for all $x\in X$. Therefore, under the correspondence $\rho_{e,f}$, the underlying group $(X,\circ_{e,f})$ can be identified with the displacement group and without loss of generality we can say that $X$ is right linear over $\mathrm{Dis}(X)$.
By the main theorem of [16] and 26, we have that the orbits of a minimal normal subgroup of $\mathcal{G}(X)$ forms a non-trivial complete blocks system. 30 and 31 allow to describe all the complete blocks systems and all the congruences of irretractable cycle sets with regular displacement group in terms of normal subgroups and ideals contained in $\mathrm{Dis}(X)$.
Corollary 34.
Let $X$ be an irretractable cycle set with regular displacement group and let $e,f$ be elements of $X$. Then:
  • 1)
    the complete blocks systems are in one-to-one correspondence with the subgroups of $\mathrm{Dis}(X)$ that are invariant by $\overline{\psi_{e,f}}$;
  • 2)
    the congruence of $X$ are in one-to-one correspondence with normal subgroups $H$ of $\mathrm{Dis}(X)$ that are invariant under $\overline{\psi_{e,f}}$ and $\overline{\phi_{e,f}}$.
Proof.
1)1)1 ) and 2)2)2 ) follows by 33, 30, 31, and 2. Since the displacement group $\mathrm{Dis}(X)$ is the set of left translations of $(X,\circ_{e,f})$, the last part of the statement also follows. ∎
The congruences of $X$, when it is an irretractable cycle set with regular displacement group, can be described in terms of the left brace $\mathcal{G}(X)$.
Theorem 35.
Let $X$ be a finite irretractable cycle set with regular displacement group $\mathrm{Dis}(X)$. Then, the congruences of $X$ are in one-to-one correspondence with ideals of $\mathcal{G}(X)$ contained in $\mathrm{Dis}(X)$.
Proof.
By 33, $X$ can be identified with a cycle set right linear over a group $(G,\circ)$ and given by $x\cdotp y:=\phi(x)\circ\psi(y)$ for all $x,y\in G$, for a suitable maps $\phi$ and $\psi$. By 31 the congruences of $X$ are in one-to-one correspondence with normal subgroups $H$ of $G$ that are invariant under $\phi$ and $\psi$. Since $\mathrm{Dis}(X)$ acts on $X$ by left translations (respect to the operation $\circ$), the equivalence classes of the congruence induced by a subgroup $H$ of $G$ correspond to the orbits respect to the action of the subgroup $T_{H}$ of $\mathrm{Dis}(X)$ given by the translations by the elements of $H$. Therefore, since $X$ is irretractable, by [3, Lemma 2.1] and [35, Theorem 1] $T_{H}$ is an ideal of $\mathcal{G}(X)$. Therefore, every congruence comes from a suitable ideal of $\mathcal{G}(X)$ contained in $\mathrm{Dis}(X)$. Since $\mathrm{Dis}(X)$ acts by left translation with respect to the operation $\circ$, different ideals contained in $\mathrm{Dis}(X)$ give rise to different congruences, hence the statement follows. ∎
The following result describes the left brace structure of the permutation group $\mathcal{G}(X)$ of an irretractable cycle set with regular displacement group of prime-power size.
Proposition 36.
Let $X$ be a finite irretractable cycle set with regular displacement group and $e,f$ be elements of $X$. Then, the permutation group $\mathcal{G}(X)$ is isomorphic to $\mathrm{Dis}(X)\rtimes<\psi_{e,f}>$.
Proof.
It follows by 33 and 2, toegether with a standard calculation. ∎
Corollary 37.
Let $p$ be a prime. Let $X$ be an irretractable cycle set with regular displacement group of size $p^{k}$. Then, $\mathcal{G}(X)$, as a left brace, is isomorphic to the semidirect product of $B_{p}$ and $C_{n}$, where $B_{p}$ is the $p$-Sylow subgroup of $(\mathcal{G}(X),+)$ and $C_{n}$ is a trivial cyclic left brace and $gcd(n,p)=1$.
Proof.
It follows by 33, 36 and 23. ∎
As a final remark of this section, note that, since an abelian transitive subgroup is regular, an irretractabe cycle set $X$ with abelian displacement group $\mathrm{Dis}(X)$ is latin if and only if $\mathrm{Dis}(X)$ acts transitively on $X$. Therefore, our theory applies to all cycle sets with regular and abelian displacement group.
5 Simple cycle sets
5.1 Simple left quasigroups
Let us turn our attention to simple left quasigroups.
Definition 7.
Let $X$ be an indecomposable left quasigroup. Then, $X$ is simple if the only congruences of $X$ are $0_{X}$ and $1_{X}$.
Corollary 38.
Let $X$ be an indecomposable and irretractable left quasigroup. Then, $X$ is simple if and only if $\mathrm{Dis}(X)$ is the smallest nonzero admissible subgroup. Moreover, $\mathrm{Dis}(X)$ acts transitively on $X$.
Proof.
Suppose that $X$ is simple. Clearly, $\mathrm{Dis}(X)$ is a nonzero admissible subgroup and by the theory after 1 it acts transitively on $X$. For the same reason, every admissible subgroup $H$ must acts transitively on $X$ and, by the definition of admissible subgroup, $\mathrm{Dis}(X)$ is contained in $H$, hence $\mathrm{Dis}(X)$ is the smallest nonzero admissible subgroup of $\mathcal{G}(X)$.
Conversely, suppose that $\mathrm{Dis}(X)$ is the smallest nonzero admissible subgroup that acts transitively on $X$. Then, this implies the transitivity of $H$ on $X$, hence simplicity of $X$ follows by 6. ∎
From now on, a cycle set will be called simple if its underlying left quasigroup is simple. As it is well-known that if $X$ is an indecomposable cycle set then the ideal $\mathcal{G}(X)^{2}$ of $\mathcal{G}(X)$ coincides with $\mathrm{Dis}(X)$ (see [11, Lemma 11]), we can use our results to recover a characterization of simple cycle sets. Since the unique simple decomposable cycle set has size $2$, we decide to add the condition $|X|>2$ and isolate this easy case.
Corollary 39.
Let $X$ be a finite cycle set and suppose that $|X|>2$. Then, $X$ is simple if and only if $X$ is the indecomposable trivial cycle set of prime size $p$ or $X$ is irretractable and $\mathcal{G}(X)^{2}$ is a minimal ideal of $\mathcal{G}(X)$ that acts transitively on $X$.
Proof.
If $X$ is simple and is not the trivial cycle set of prime size, by [17, Proposition 4.1] it is irretractable, hence this implication follows by 20, 38 and the equality $\mathcal{G}(X)^{2}=\mathrm{Dis}(X)$. Conversely, if $X$ is the trivial indecomposable of prime size, then it is simple by [12, Lemma 1]. Now, suppose that $X$ is irretractable and $\mathcal{G}(X)^{2}$ is a minimal ideal of $\mathcal{G}(X)$ that acts transitively on $X$. Then $X$ is indecomposable and hence we have $\mathcal{G}(X)^{2}=\mathrm{Dis}(X)$. Therefore, the thesis follows by 20 and 38. ∎
5.2 Simple cycle sets with nilpotent displacement group
In this section we provide some limitations on the cardinality of simple cycle sets having nilpotent permutation group. In particular, we show that simple cycle sets with cyclic displacement group admit a very simple description.
Lemma 40.
Let $X$ be an indecomposable cycle set and assume that $\mathrm{Dis}(X)$ has a non-trivial normal Hall subgroup $H$. Then, $H$ is an ideal of $\mathcal{G}(X)$.
Proof.
If $H$ is a non-trivial normal Hall subgroup of $\mathrm{Dis}(X)$, then by [4, Proposition 4.2.1] it is an ideal of $\mathrm{Dis}(X)$. Moreover, $H$ is a characteristic subgroup of $(\mathrm{Dis}(X),+)$ and $(\mathrm{Dis}(X),\circ)$, therefore it is an ideal of $\mathcal{G}(X)$. ∎
Corollary 41.
Let $X$ be a simple cycle set with nilpotent displacement group. Then, $|X|=p^{k}$ for a prime number $p$ and a natural number $k$.
Proof.
If $p,q$ are distinct prime numbers dividing $|X|$, then the $q$-Sylow of $(\mathrm{Dis}(X),\circ)$ is a non-trivial normal Hall subgroup of $(\mathrm{Dis}(X),\circ)$. By Lemma 40 and 21 $X$ has a non-trivial congruence, a contradiction. ∎
Now, we can describe all the simple cycle sets with cyclic displacement group.
Theorem 42.
Let $X$ be a finite simple cycle set. Then, the following statements are equivalent:
  • 1)
    $\mathrm{Dis}(X)$ is cyclic;
  • 2)
    $|\mathrm{Dis}(X)|=1$;
  • 3)
    $|X|=2$ or $|X|$ is an indecomposable cycle set of prime size.
Proof.
1) $\Rightarrow$ 3) Suppose that $X$ is a simple cycle set with cyclic displacement group and such that $|X|$ is not a prime number. Then, $X$ must be irretractable and hence $\mathrm{Dis}(X)$ acts transitively on $X$, therefore by 41 we have $|X|=|\mathrm{Dis}(X)|=p^{k}$ for some prime number $p$. Moreover, by an inspection of the small cycle sets, $|X|\neq 4$. Now, let $m$ be a divisor of $|X|$ such that $|X|/m=p$ for a prime number $p$. By the results containted in [36] and [2, Proposition 5.4], $\mathrm{Dis}(X)$ is a left brace with cyclic additive and multiplicative group. Moreover, the set $H:=p\mathrm{Dis}(X)$ is an ideal of $\mathrm{Dis}(X)$ of size $m$, and is a characteristic subgroup of $(\mathrm{Dis}(X),+)$ and $(\mathrm{Dis}(X),\circ)$, hence it is an ideal of $\mathcal{G}(X)$. Therefore, $p\mathrm{Dis}(X)$ induces a non-trivial congruence on $X$, a contradiction. Then, $|X|$ must be a prime number, and by [17, Proposition 4.1] $|X|=2$ or $X$ is indecomposable. If the first case, clearly $\mathrm{Dis}(X)=\{id_{X}\}$ by a standard calculation. In the second case the conclusion is the same by [25, Theorem 2.13].
3) $\Rightarrow$ 2) If $X$ is a simple cycle set of prime size, we have that $\mathrm{Dis}(X)$ is the trivial group by [25, Theorem 2.13]
2) $\Rightarrow$ 1) Clear. ∎
5.3 Simple cycle sets with regular displacement group
In this subsection, we provide a description of simple cycle sets with nilpotent regular permutation group.
Corollary 43.
Let $X$ be a finite simple cycle set with regular displacement group. Then, $X$ is right linear over $\mathrm{Dis}(X)$.
Proof.
The simplicity of $X$ implies that $X$ is irretractable or $|\mathrm{Dis}(X)|=1$. Since $\mathrm{Dis}(X)$ acts regularly, $X$ must be irretractable, hence the statement follows by 33. ∎
From now on, in this subsection we indicate by $(p,G,\phi,\psi)$ a $4$-uple such that:
  • 1.
    $p$ is a prime number
  • 2.
    $(G,+)$ is a group of size $p^{k}$;
  • 3.
    $(\phi,\psi)\in Sym(G)\times Aut(G,+)$ such that Eq. 8 holds for all $x,y\in G$.
By 34, 41 and 43, we are able to construct all the simple cycle sets with nilpotent regular permutation group. We leave the proof to the reader since it follows by the cited results.
Theorem 44.
Let $(p,G,\phi,\psi)$ be such that every normal subgroup $H$ of $G$ invariant under $\psi$ is not invariant under $\phi$. Then, the pair $(X,\bullet)$ given by $X:=G$ and
$x\bullet y:=\phi(x)+\psi(y)$
for all $x,y\in G$ is a non-trivial simple cycle set with $\mathrm{Dis}(X)\cong G$.
Conversely, every simple cycle set $X$ having nilpotent regular displacement group can be constructed in this way.
If $X$ is a simple cycle set with nilpotent regular displacement group, by 22 it can be constructed by a suitable cycle set having a $p$-group permutation group.
Corollary 45.
Let $Y$ be a simple cycle set having nilpotent regular displacement group. Then, it is a deformation of a right linear cycle set $X$ by a suitable $\alpha\in Aut(X)$ such that $\mathcal{G}(X)$ is a $p$-group for some prime number $p$.
Proof.
It follows by 41, 43, 37, 22, and 29. ∎
By the previous corollary, if $Y$ is a simple cycle set with nilpotent regular displacement group, then it is a deformation of some cycle set $X$ by an automorphism $\alpha$. Unfortunately, we do not know if such an $X$ is simple. Moreover, we do not know if the deformation of a simple cycle set is again simple, but we are able to show this under some additional hypothesis.
Proposition 46.
Let $X$ be a simple cycle set having a $p$-group permutation group $\mathcal{G}(X)$. Let $\alpha$ be an automorphism of $X$ having order coprime with $p$ and fixing an element $x\in X$. Then, the deformation $X_{\alpha}$ of $X$ is simple.
Proof.
Suppose that $X_{\alpha}$ is not a simple cycle set. Then, by [22, Lemma 4.5] the left brace $\mathcal{G}(X_{\alpha})$ is isomorphic to $\mathcal{G}(X)\rtimes<\alpha>$, and by 39 there exist a nonzero ideal $J$ of $\mathcal{G}(X_{\alpha})$ that acts not transitively on $X_{\alpha}$. Then, the ideal $J\cap\mathrm{Dis}(X_{\alpha})$, which is equal to $J\cap\mathcal{G}(X_{\alpha})^{2}$, contains $J*\mathcal{G}(X)$, therefore if $J\cap\mathrm{Dis}(X_{\alpha})=0$ we have $J*\mathcal{G}(X_{\alpha})=0$ and hence $J\subseteq Soc(\mathcal{G}(X_{\alpha}))=0$, but this is not possible because $J$ is nonzero. Hence $J\cap\mathrm{Dis}(X_{\alpha})$ is a nonzero ideal of $\mathcal{G}(X_{\alpha})$ contained in $\mathrm{Dis}(X_{\alpha})$. Since by 29 $\mathrm{Dis}(X_{\alpha})=\mathrm{Dis}(X)$, we have that $J\cap\mathrm{Dis}(X)$ is a nonzero ideal of $\mathcal{G}(X)$ that acts not transitively on $X$, hence by 39 $X$ is not simple, a contradiction. ∎
Corollary 47.
Let $(p,G,\phi,\psi)$ be such that $\psi$ has order $p^{s}$ for some natural number $s$, and every normal subgroup $H$ of $G$ invariant under $\psi$ is not invariant under $\phi$. Moreover, let $\alpha\in Aut(G,\cdotp)$, where $x\cdotp y:=\phi(x)+\psi(y)$ for all $x,y\in G$, such that $\alpha$ has coprime order with $p$ and $\alpha(z)=z$ for some $z\in G$. Then, the pair $(X,\bullet)$ given by $X:=G$ and
$x\bullet y:=\alpha(\phi(x)+\psi(y))$
for all $x,y\in G$ is a simple cycle set.
Proof.
It follows by 37, 44 and 46. ∎
6 Affine cycle sets
[…]
$x\cdotp y:=\phi(x)+\psi(y)+c$ (9)
for all $x,y\in X$, where $\phi,\psi\in Aut(X,+)$ and $c\in X$. Recall that we denote such a cycle set by $\mathrm{Aff}(X,\phi,\psi,c)$
The following result, that is a reformulation of [9, Theorem $4.18$], allows to detect the class of affine cycle sets by group-theoretic properties.
Corollary 48.
Let $X$ be a finite irretractable cycle set with abelian displacement group $\mathrm{Dis}(X)$. Then, $X$ is affine if and only if $\mathrm{Dis}(X)$ is a normal subgroup of $\mathcal{TG}(X)$ that acts regularly on $X$.
Proof.
It follows by 32 and [9, Theorem $4.18$]. ∎
Latin affine cycle sets can be understood as modules over the Weyl algebra over the integers.
Remark 3.
Let us define the first Weyl algebra over a commutative ring $R$ with unit $1$ as $A_{1}(R)=R[a,b]/\langle ab-ba-1\rangle$.
  1. 1.
    Let $(X,\cdotp)=\mathrm{Aff}(A,\phi,\psi,c)$ be a latin affine cycle set with underlying abelian group $(A,+)$. As showed in [9, Section $4$], for affine cycle sets Eq. 6 is equivalent to the relation $[\psi,\phi^{-1}]=id_{X}$, that is also the defining relation of the first Weyl Algebra $A_{1}(\mathbb{Z})$. Thus $(X,+)$ has a structure of $A_{1}(\mathbb{Z})$-module defined by $\rho:A_{1}(\mathbb{Z})\longrightarrow\operatorname{End}{(A,+)}$ with $\rho(a)=\psi$ and $\rho(b)=\phi^{-1}$.
  2. 2.
    Conversely, if $M$ is a $A_{1}(\mathbb{Z})$-module, defined by $\rho:A_{1}(\mathbb{Z})\rightarrow\operatorname{End}{(M,+)}$ with $\rho(a),\rho(b)\in\operatorname{Aut}{(M,+)}$ then $\mathrm{Aff}(M,\rho(b)^{-1},\rho(a),c)$ gives rise to an affine cycle set for every $c\in M$.
Note that, if $(M,+)$ is an elementary abelian $p$-group, then $(M,+)$ is a $\mathbb{Z}/p\mathbb{Z}$-module and so $\mathrm{Aff}(M,\phi,\psi,c)$ is a $A_{1}(\mathbb{Z}/p\mathbb{Z})$-module. Hence we have a representation of the algebra $A_{1}(\mathbb{Z}/p\mathbb{Z})$ over $M$, and so the dimension of the representation must be a multiple of $p$ [9, Proposition 4.7].
6.1 Simple affine cycle sets
The properties of being affine and simple force the displacement group to be elementary abelian.
Proposition 49.
Let $X$ be a simple affine latin cycle set. Then, $\mathrm{Dis}(X)$ is an elementary abelian $p$-group of size $p^{k}$ for some prime number $p$ dividing $k$.
Proof.
Since $X$ is affine then $\mathrm{Dis}(X)$ is abelian. By 41 the underlying abelian group is a p-group. Since in a finite abelian group $(G,+)$ the elements of the form $pg$, with $g\in G$, form a characteristic subgroup, the statement follows by 34 and 2. ∎
Let $p$ be a prime number. As a main consequence of 3 and 49, we obtain that every finite simple affine cycle set corresponds to a finite-dimensional irreducible representations $\rho$ of $A_{1}(\mathbb{Z}/p\mathbb{Z})$ together with a constant $c\in(\mathbb{Z}/p\mathbb{Z})^{n}$, where $n:=dim(\rho)$.
Theorem 50.
Let $p$ be a prime number and $\rho$ be an irreducible representation with dimension $n$ of $A_{1}(\mathbb{Z}/p\mathbb{Z})$. Suppose that $\rho(a)$ and $\rho(b)$ are invertible and set $\phi:=\rho(b)^{-1}$ and $\psi:=\rho(a)$. Then $\mathrm{Aff}(\mathbb{Z}/p\mathbb{Z})^{n},\phi,\psi,c)$is a simple latin affine cycle set for every $c\in\mathbb{Z}/p\mathbb{Z}$. Moreover, $\mathcal{G}(X)$ is a $p$-group if and only if $\psi$ has prime-power order.
Conversely, every affine simple latin cycle set can be constructed in this way.
Proof.
[…]
Conversely, suppose that $X$ is an affine simple latin cycle set. Then, by [9, Section 4] and 49 there exist a natural number $n$, a prime number $p$ dividing $n$, an element $c\in(\mathbb{Z}/p\mathbb{Z})^{n}$, and two automorphism $\phi$ and $\psi$ of $(\mathbb{Z}/p\mathbb{Z})^{n}$, such that $X:=(\mathbb{Z}/p\mathbb{Z})^{n}$ and $x\cdotp y:=\phi(x)+\psi(y)+c$ for all $x,y\in(\mathbb{Z}/p\mathbb{Z})^{n}$. Then, the assignments $a\mapsto\psi$ and $b\mapsto\phi^{-1}$ gives rise to a representation $\rho$ of $A_{1}(\mathbb{Z}/p\mathbb{Z})$ (see also Remark 3). By simplicity of $X$, the irreducibility of $\rho$ follows. ∎
Remark 4.
We highlight that by 48, the simple cycle sets $X$ with displacement group abelian and normal in the total permutation group $\mathcal{TG}(X)$ are all and only the ones provided in the previous theorem.
Now, we indicate by $(p,\rho,c,\eta,g)$ a quintuple such that:
  • 1.
    $p$ is a prime number;
  • 2.
    $\rho$ is a representation of the Weyl Algebra $A_{1}(\mathbb{Z}/p\mathbb{Z})$ with $\rho(a):=\psi$ and $\rho(b):=\phi$ invertible matrices;
  • 3.
    $c\in(\mathbb{Z}/p\mathbb{Z})^{n}$, where $n:=dim(\rho)$;
  • 4.
    $(\eta,g)\in C_{Aut(G,+)}(A,B)\times(\mathbb{Z}/p\mathbb{Z})^{n}$ is such that $\eta(c)-c=Ag+Bg-g$.
Below, we show an affine-version of 47. The proof is left to the reader, since it follows by the previous results and is standard.
Corollary 51.
Let $(p,\rho,c,\eta,g)$ be a quintuple such that $\rho$ is an irreducible representation, $\psi$ as order $p^{s}$ for some natural number $s$, and $\eta$ has order coprime with $p$. Then, the pair $((\mathbb{Z}/p\mathbb{Z})^{n},\cdotp)$ with
$x\cdotp y:=g+\eta(\rho(b)^{-1}x+\rho(a)y+c)$
for all $x,y\in(\mathbb{Z}/p\mathbb{Z})^{n}$ is a simple cycle set having $\mathrm{Dis}(X)$ abelian, regular and normal in $\mathcal{TG}(X)$.
We close this subsection showing that even if, in full generality, an affine cycle set is not simple (for example, take an affine cycle set having size $p^{\alpha}q^{\beta}$, for distinct prime numbers $p,q$), it always provides simple non-trivial cycle sets.
Proposition 52.
Let $X$ be a finite cycle set affine over a group $G$ and given by $x\cdotp y:=\phi(x)+\psi(y)+c$. Let $H$ be a subgroup of $G$ that is minimal among the subgroups invariant under the actions of $\phi$ and $\psi$. Let $h\in H$ and define $Y$ as the cycle set given by $Y:=H$ and $x\cdotp y:=\phi(x)+\psi(y)+h$ for all $x,y\in Y$. Then, $Y$ is a simple non-trivial cycle set affine over $H$.
Proof.
By a standard calculation, one can show that $Y$ is a non-trivial cycle set affine over $H$. If $Y$ is not simple, there exist a subgroup $K$ of $H$ that is invariant under the actions of $\phi$ and $\psi$. This implies that $H$ is not minimal among the subgroups of $G$ invariant under the actions of $\phi$ and $\psi$, a contradiction. Hence $Y$ is simple. ∎
6.2 Simple affine cycle sets of minimal size
By 50, several examples of simple cycle sets can be obtained by irreducible representations of the first Weyl Algebra $A_{1}(\mathbb{Z}/p\mathbb{Z})$. Even if these representations are not classified over arbitrary fields (to our knowledge), several examples can be recovered from the irreducible representations of the first Weyl Algebras on algebrically closed fields, that are completely classified. Indeed, if $K$ is an algebrically closed field of characteristic $p>0$, all the finite dimensional irreducible representations of $A_{1}(K):=K[a,b]/(ab-ba-1)$ are given by the $p\times p$ matrices
$M_{a}:=\begin{bmatrix}0&0&0&...&\mu\\ 1&0&0&...&0\\ 0&1&0&...&0\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&0&1&0\end{bmatrix}\qquad and\qquad M_{b}:=\begin{bmatrix}\lambda&p-1&0&...&0\\ 0&\lambda&p-2&...&0\\ 0&0&\lambda&...&0\\ \vdots&\vdots&\vdots&\vdots&1\\ 0&0&0&0&\lambda\end{bmatrix}$ (10)
where $\mu,\lambda\in K$. If we restrict to representations of dimension $p$, all the irreducible ones of $A_{1}(K)$ are given by a pair of matrices as above. The argument for the proof of the following Proposition was suggested in [28].
Proposition 53.
Let $p$ be a prime number, $V:=(\mathbb{Z}/p\mathbb{Z})^{p}$ and $\rho:A_{1}(\mathbb{Z}/p\mathbb{Z})\rightarrow End(V)$ be an irreducible representation of $A_{1}(\mathbb{Z}/p\mathbb{Z})$ of dimension $p$. Then, there exists a basis $\{e_{1},...,e_{p}\}$ of $V$ such that $\rho(a)=M_{a}$ and $\rho(b)=M_{b}$ for suitable elements $\mu,\lambda\in\mathbb{Z}/p\mathbb{Z}$.
Proof.
Let $M_{1}:=\rho(a)$, $M_{2}:=\rho(b)$ with respect some basis =$\mathcal{B}=\{e_{1},\ldots,e_{p}\}$ of $V$, and $K$ be the algebraic closure of $\mathbb{Z}/p\mathbb{Z}$. Since $dim(V)=p$, $M_{1}$ and $M_{2}$ provide a $p$-dimensional representation $\bar{\rho}$ of $A_{1}(K)$, that must be irreducible. Moreover, we have that $\mathcal{B}$ can be regarded as a basis of $K^{p}$, and we have $\bar{\rho}(a)=M_{1}$ and $\bar{\rho}(b)=M_{2}$ with respect $\mathcal{B}$. By the previous remark, there exist a basis :=$\mathcal{B}^{\prime}:=\{e^{\prime}_{1},...,e^{\prime}_{p}\}$ of $K^{p}$ such that $\bar{\rho}(a)=M_{a}$ and $\bar{\rho}(b)=M_{b}$ for suitable elements $\mu,\lambda\in K$. Hence, there exist a matrix $C\in GL_{p}(K)$ providing the change of basis from $\mathcal{B}^{\prime}$ to $\mathcal{B}$ such that $M_{a}=C^{-1}M_{1}C$ and $M_{b}=C^{-1}M_{2}C$. To show the statement, is sufficient proving that $C\in GL_{p}(\mathbb{Z}/p\mathbb{Z})$.
The characteristic polynomial does not change by conjugacy, therefore we have that the characteristic polynomial of $M_{2}$, which is equal to the one of $M_{b}$, is $(x-\lambda)^{p}=x^{p}-\lambda^{p}$, hence $\lambda^{p}\in\mathbb{Z}/p\mathbb{Z}$. If $F$ is the Frobenius endomorphism of $\mathbb{Z}/p\mathbb{Z}$, there exist $\alpha\in\mathbb{Z}/p\mathbb{Z}$ such that $F(\alpha)=\lambda^{p}$, hence $\alpha^{p}=\lambda^{p}$, i.e. $\alpha$ is a solution of the polynomial $(x-\lambda)^{p}$. This proves $\alpha=\lambda$. Since the dimension of the $\lambda$-eigenspace of $M_{2}$ (which is equal to the dimension of the $\lambda$-eigenspace of $M_{b}$) is $1$, the vector $e^{\prime}_{1}$, up to multiply by a suitable scalar, is an eigenvector of $M_{2}$. Therefore, we can assume that $e^{\prime}_{1}=e_{1}$ (up to consider matrices $M_{1}^{\prime}$ and $M_{2}^{\prime}$ similar to $M_{1}$ and $M_{2}$ in $GL_{p}(\mathbb{Z}/p\mathbb{Z})$). Now, let $f_{i}\in\mathbb{Z}/p\mathbb{Z}^{p}$ having 00 in the $j$-th position, with $j\neq i$, and $1$ in the $i$-th position. Therefore, the first column of $C$ is $f_{1}$. Now, we have that $M_{a}^{j-1}f_{1}=C^{-1}M_{1}^{j-1}Cf_{1}=C^{-1}M_{1}^{j-1}f_{1}$ and hence $Cf_{j}=M_{1}^{j-1}f_{1}$ for every $j\in\{1,...,p\}$, and in a similar way $\mu Cf_{1}=M_{1}^{p}f_{1}$. From these equalities, we obtain $\mu\in\mathbb{Z}/p\mathbb{Z}$ and $C\in GL_{p}(\mathbb{Z}/p\mathbb{Z})$.
By 50 and 53, we are able to construct concretely all the affine simple cycle sets of size $p^{p}$, for an arbitrary prime number $p$. The next goal is to distinguish the isomorphism classes.
Now, we define the algebra of the $\mu$-circulant matrices over a field $K$ in analogy with the classical circulant matrices. Recall that a square matrix is circulant if it is constant on all broken diagonals, and denote by $C(c_{1},\ldots,c_{n})$ the $n\times n$ circulant matrix with first row equal to $(c_{1},\ldots,c_{n})$. Let $\mu\in K$ we define $C_{\mu}(c_{1},\ldots,c_{n})$ to be the matrix obtained by $C(c_{1},\ldots,c_{n})$ multiplying all the upper-triangular entries of $C$ by the element $\mu$, as
$C_{\mu}(c_{1},\ldots,c_{n})=\begin{bmatrix}c_{1}&\mu c_{2}&\ldots&\ldots&\mu c_{n-1}&\mu c_{n}\\ c_{n}&c_{1}&\mu c_{2}&\ldots&\ldots&\mu c_{n-1}\\ c_{n-1}&c_{n}&c_{1}&\mu c_{2}&\ldots&\ldots\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ c_{3}&\ldots&\ldots&c_{n}&c_{1}&\mu c_{2}\\ c_{2}&c_{3}&\ldots&c_{n-1}&c_{n}&c_{1}\end{bmatrix}\,.$ (11)
The algebra of the circulant matrices coincides with the centralizer of the permutation matrix representing the $n$-cycle. We will show that the $\lambda$-circulant matrices are the centralizer of the following matrix:
$A_{\mu}=\begin{bmatrix}0&0&\ldots&0&\mu\\ 1&0&\ldots&0&0\\ 0&1&\ldots&0&0\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\ldots&1&0\\ \end{bmatrix}.$ (12)
Lemma 54.
The set of the $n\times n$ $\mu$-circulant matrices is the subalgebra of the matrices which commutes with $A_{\mu}$ (as defined in (12)).
Proof.
Let $C=(c_{i,j})$ be a $n\times n$ matrix. Then
$(CA_{\mu})_{i,j}=\begin{cases}c_{i,j+1},\quad\text{if }j\neq n,\\ \mu c_{i,1},\quad\text{if }j=n,\end{cases},\quad(A_{\mu}C)_{i,j}=\begin{cases}c_{i-1,j},\quad\text{if }i\neq 1,\\ \mu c_{n,j},\quad\text{if }i=1.\end{cases}$
Therefore $C$ commutes with $A_{\mu}$ if and only if
$\begin{cases}c_{i,j+1}=c_{i-1,j},\,\text{ for every }i\neq 1,\,j\neq n,\\ c_{i,1}=c_{i-1,n},\,\text{ for every }i\neq 1,\\ c_{1,j+1}=\mu c_{n,j},\,\text{ for every }j\neq n,\\ \mu c_{1,1}=\mu c_{n,n}.\end{cases}$
Hence, the centralizer of $A_{\mu}$ coincides with the set of the $\mu$-circulant matrices. ∎
Let $\Delta=(d_{i,j})$ be the $n\times n$ matrix defined by
$d_{i,j}=\begin{cases}n-i,\,\text{ if }j=i+1\pmod{n},\\ 0,\,\text{otherwise}.\end{cases}$
e.g. if $n=3$ we have
$\Delta=\begin{bmatrix}0&2&0\\ 0&0&1\\ 0&0&0\end{bmatrix}$
Lemma 55.
Let $C$ be a $p\times p$ $\mu$-circulant matrix. Then $[C,\Delta]=0$ if and only if $C=sI$ for some $s\in K$.
Proof.
Let $C=\mu$-$Circ(c_{1},\ldots,c_{p})$. It is straightforward to compute that the first column of $\Delta\cdotp C$ is $((p-1)c_{p},(p-2)c_{p-1},\ldots,c_{2},0)$ and the first column $C\cdotp\Delta$ is $(0,\ldots,0)$. Therefore if $[C,\Delta]=0$ holds then $c_{2}=c_{3}=\ldots=c_{p}=0$ and so $C=c_{1}I$. ∎
To identify the isomorphism classes of affine cycle sets of size $p^{p}$ we are going to use the following criterion. The criterion is state for quasigroups and so it covers the case of finite latin cycle sets.
Theorem 56 (Drápal [23, Thm. 3.2]).
Let $Q=\mathrm{Aff}(G,\phi,\psi,c)$ and $Q^{\prime}=\mathrm{Aff}(G,\phi^{\prime},\psi^{\prime},c^{\prime})$ be affine quasigroups. Then $Q$ is isomorphic to $Q^{\prime}$ if and only if there are $\alpha\in\operatorname{Aut}{(G,+)}$ and $u\in\mathrm{Im}(1-phi-psi)$ such that $\phi^{\prime}=\alpha\phi\alpha^{-1}$, $\psi^{\prime}=\alpha\psi\alpha^{-1}$ and $c^{\prime}=\alpha(c+u)$.
Theorem 57.
Let $p$ be a prime and $X$ be an affine simple latin cycle set of size $p^{p}$. Then $X$ is isomorphic to one of the following latin cycle sets:
$\mathrm{Aff}(\mathbb{Z}_{p}^{p},(\lambda I+\Delta)^{-1},\mathcal{A}_{\mu},(0,\ldots,0)),\quad\quad\mathrm{Aff}(\mathbb{Z}_{p}^{p},(\lambda I+\Delta)^{-1},\mathcal{A}_{1},(1,0,\ldots,0)).$
for $\mu,\lambda=1,\ldots,p-1$. In particular, there are $p(p-1)$ affine simple latin cycle sets of size $p^{p}$.
Proof.
According to 50 and 53, we have that $X$ is given by $X(\lambda,\mu,c)=\mathrm{Aff}(\mathbb{Z}_{p}^{p},(\lambda I+\Delta)^{-1},\mathcal{A}_{\mu},c)$. If $X(\lambda,\mu,c)$ and $X(\lambda^{\prime},\mu^{\prime},c^{\prime})$ are isomorphic then the matrices $A_{\mu}$ and $A_{\mu^{\prime}}$ (resp. $\mu I+\Delta$ and $\mu^{\prime}I+\Delta$) are conjugate according to 56. Therefore such matrices have the same eigenvalues, i.e. $\lambda=\lambda^{\prime}$ (resp. $\mu=\mu^{\prime}$). By this fact and 56, $X(\lambda,\mu,c)$ and $X(\lambda^{\prime},\mu^{\prime},c^{\prime})$ are isomorphic if and only if and only if $\lambda=\lambda^{\prime}$, $\mu=\mu^{\prime}$ and there exists $\alpha\in GL_{p}(\mathbb{Z}/p\mathbb{Z})$ and $u\in\mathrm{Im}(1-(\lambda I+\Delta)^{-1}-\mathcal{A}_{\mu})$ such that $\mu I+\Delta=\alpha(\mu I+\Delta)\alpha^{-1}$, $\mathcal{A}_{\mu}=\alpha\mathcal{A}_{\mu}\alpha^{-1}$ and $c^{\prime}=\alpha(c+u)$. By virtue of Lemma 55 then $\alpha=sI$ with $s\neq 0$. Therefore
$c^{\prime}-sc=su\in\mathrm{Im}(1-(\lambda I+\Delta)^{-1}-\mathcal{A}_{\mu}),$
i.e. $c$ and $c^{\prime}$ are generate the same subspace modulo $\mathrm{Im}(1-(\lambda I+\Delta)^{-1}-\mathcal{A}_{\mu})$. Thus, we need to understand the dimension of $\mathrm{Im}(1-(\lambda I+\Delta)^{-1}-\mathcal{A}_{\mu})$ in order to account how many different isomorphism classes we have. The dimension of $\mathrm{Im}(1-(\lambda I+\Delta)^{-1}-\mathcal{A}_{\mu})$ equals the dimension of
$\mathrm{Im}(\lambda I+\Delta)(1-(\lambda I+\Delta)^{-1}-\mathcal{A}_{\mu})=\mathrm{Im}((\lambda I+\Delta)(1-\mathcal{A}_{\mu})-1).$
Now, we have that $(\lambda I+\Delta)(1-\mathcal{A}_{\mu})-1$ can be written as
$\begin{bmatrix}\lambda&p-1&0&0&0&\ldots&-\lambda\mu\\ -\lambda&\lambda+1&p-2&0&0&\ldots&0\\ 0&-\lambda&\lambda+2&p-3&0&\ldots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ \vdots&\vdots&0&-\lambda&\lambda+p-3&2&0\\ 0&\ldots&0&0&-\lambda&\lambda+p-2&1\\ 0&&\ldots&0&0&-\lambda&\lambda+p-1\end{bmatrix}.$
By a standard linear algebra argument, we have that the rank of the matrix above is $p$ if $\mu\neq 1$ and $p-1$ if $\mu=1$. Thus, in the first case the rank of $\mathrm{Im}(1-(\lambda I+\Delta)^{-1}-A_{\mu})$ is $p$ and hence two arbitrary constant $c,c^{\prime}$ provide isomorphic solutions. Now, let $\mu=1$ and so the rank of $\mathrm{Im}((\lambda I+\Delta)(1-\mathcal{A}_{1})-1)$ is $p-1$. By a standard linear algebra argument, we have that $\mathrm{Im}((\lambda I+\Delta)(1-\mathcal{A}_{1})-1)$ is the subspace given by the vectors $(x_{1},\ldots,x_{p})$ with $x_{1}+...+x_{p}=0$ and therefore the vector $(1,0,\ldots,0)$ does not belong to this space. Moreover, the vector $(1,0,\ldots,0)$ does not belong to $\mathrm{Im}(1-(\lambda I+\Delta)^{-1}-\mathcal{A}_{\mu})$, because otherwise, being $(1,0,\ldots,0)$ an eigenvector of $(\lambda I+\Delta)$, it would imply that $(1,0,\ldots,0)\in\mathrm{Im}((\lambda I+\Delta)(1-(\lambda I+\Delta)^{-1}-\mathcal{A}_{\mu}))$, a contradiction. By a standard calculation, we obtain that there are two isomorphism classes of simple cycle sets of the form $X(\lambda,1,c)$, one is given by $c:=(0,\ldots,0)$ and the other one is given by $c:=(1,0,\ldots,0)$. Hence we obtain $(p-1)\cdotp(p-2)$ simple affine cycle sets for $\mu\neq 1$ and $2\cdotp(p-1)$ for $\mu=1$, therefore the statement follows. ∎
7 Applications and examples
In this section, we collect examples of simple latin cycle sets obtained by our results and we discuss some applications to open questions present in literature.
We start by a non-existence result that extend [9, Lemma 4.13]. In particular, we show that we have no latin cycle sets with cyclic displacement group.
Theorem 58.
Let $X$ be a finite latin cycle set with cyclic displacement group. Then, $|X|=1$.
Proof.
Let $n:=|X|$ and suppose that $\mathrm{Dis}(X)$ is a cyclic group. At first, suppose that $n$ has an odd prime divisor $p$ and let $m$ be the positive integer such that $n/m=p$. As in 42, the set $H:=p\mathrm{Dis}(X)$ is an ideal of $\mathrm{Dis}(X)$ of size $m$ and is a characteristic subgroup of $(\mathrm{Dis}(X),+)$ and $(\mathrm{Dis}(X),\circ)$, hence it is an ideal of $\mathcal{G}(X)$. Therefore, $H$ induces a congruence $\sim$ on $X$ such that $|X/\sim|=p$, and this implies that there exist a latin cycle set of size $p$, but this contradicts [25, Theorem 2.13]. Now, we consider the case $n=2^{t}$ for some natural number $t$. The case $t=1$ is not possible by [25, Theorem 2.13]. The case $t=2$ is excluded by an inspection of small cycle sets. The case $t>2$ can be showed as in the first part by [2, Proposition 5.4]. ∎
By [22] every indecomposable and irretractable cycle set of size $p^{2}$ is simple. A complete description of the latin ones follows by [22, Theorem $5.1$], as we show in the following result.
Proposition 59.
Let $X$ be a latin cycle set of size $p^{2}$. Then, $X$ is one of the two indecomposable and irretractable cycle sets of order $4$.
Proof.
By 3, $X$ must be indecomposable and irretractable. If $p=2$, we know by [9] that the two indecomposable and irretractable cycle sets of order $4$ are latin. If $p>2$, by [22, Theorem $5.1$] $X$ is given by $\mathbb{Z}/p\mathbb{Z}\times\mathbb{Z}/p\mathbb{Z}$ and $(a,x)\cdotp(b,y)=(\alpha b+\alpha x,\alpha y+\alpha\phi(b-a))$, where $\alpha\in\mathbb{Z}/p\mathbb{Z}\setminus\{0\}$ and $\phi$ is a map from $\mathbb{Z}/p\mathbb{Z}$ to $\mathbb{Z}/p\mathbb{Z}$ with $\phi(x)=\phi(-x)$ and $\phi(\alpha x)=\alpha\phi(x)$ for all $x\in\mathbb{Z}/p\mathbb{Z}$. Let $a,b\in\mathbb{Z}/p\mathbb{Z}$ with $a\neq b$ and set $a^{\prime}:=2b-a$. Since $p>2$ we have $a\neq a^{\prime}$, but $(a,x)\cdotp(b,y)=(a^{\prime},x)\cdotp(b,y)$, a contradiction. ∎
Corollary 60.
Let $X$ be a latin cycle set of size $p^{3}$. Then, $X$ is simple.
Proof.
If $p=2$, one can show by computer calculation that there are no latin cycle sets of size $8$. If $p>2$ and $X$ is not simple, then we obtain a latin cycle set having size in $\{p,p^{2}\}$, but this contradicts [25, Theorem 2.13] and 59. ∎
In the following example, we apply 44 to show the simplicity of a right linear cycle set.
Example 61.
Using the transitive cycle bases of the left brace $SmallBraces(81,804)$ of [43], we obtain an irretractable cycle set $X$ of size $27$ having the displacement group, isomorphic to the extra-special $3$-group of size $27$, that acts transitively, and hence regularly, on $X$. By 33, this cycle set is right linear over $\mathrm{Dis}(X)$. By a long but standard calculation (helping us by GAP), one can write $X$ as a cycle set on a group $(G,\circ)$ isomorphic to $\mathrm{Dis}(X)$, where $G:=\{1,...,27\}$ with $1$ as neutral element, given by $x\cdotp y:=\phi(x)\circ\psi(y)$, where
$\phi:=(1,12,27,13,18,10,11,20,17,9,19,21,22,16,26,5,15,2,4,7,23,3,25,6,14,24)$
and
$\psi:=(2,25,16)(3,14,13)(4,8,27)(5,23,21)(6,19,18)(7,26,22)(9,20,24)(10,11,17).$
The unique $\psi$-invariant subgroups are $H_{1}:=\{1,12,15\}$ and $H_{2}:=\{1,5,9,12,15,20,21,23,24\}$. On the other hand, $\phi(1)=12$ and $\phi(15)=2$, hence $H_{1}$ and $H_{2}$ are not invariant under $\phi$. Therefore, by 44, $X$ is a simple cycle set right linear over $(G,\circ)$.
In [32], the following question was posed.
Question 1.
(Question $3.16$ of [32]) Let $X$ be a cycle set. Is it true that if some $\sigma_{x}$ contains a non-trivial cycle coprime with $|X|$, then $X$ is decomposable?
For some classes of cycle sets, the answer to the previous question is positive (see [11, 29]). In [30], Kanrar and Rump found a family of counterexamples: they are non-simple cycle sets of $2-$power size and the smaller one has size $256$. Below, we show that the construction of smaller counterexamples that are also simple cycle sets is possible. At first, we need a lemma implicitly contained in [39, Section 5].
Lemma 62.
Let $X$ be an affine cycle set given by $x\cdotp y:=\phi(x)+\psi(y)$ for all $x,y\in X$. Then, $x\cdotp y=y$ if and only if $\phi(x)=(I-\psi)y$. If $(I-\psi)$ is invertible, this is equivalent to $(I-\psi)^{-1}\phi(x)=y$ and in this case every $\sigma_{x}$ has only one fixed point.
Example 63.
Let $p$ be a prime number, with $p>2$, and $X$ be the simple affine cycle set of size $p^{p}$ provided by 57 with $\lambda=1$, $\mu\neq 1$, and $c=(0,\ldots,0).$ Since the determinant of $I-B$ is $1-\mu$, by Lemma 62 every left multiplication of $X$ has only one fixed point. Let $x$ be an element of $X$. If every non-trivial cycle of $\sigma_{x}$ is not coprime with $p$, and hence divisible by $p$, we obtain that number of the fixed points of $\sigma_{x}$ is divisible by $p$, a contradiction. Therefore, $X$ is a counterexample to 1.
In [17], the following question was posed.
Question 2.
(Question $7.6$ of [17]) Does there exist a finite simple cycle set $X$ such that $X=Y\times Z$, the sets $X_{y}=\{(y,z)\mid z\in Z\}$, for $y\in Y$, are blocks of imprimitivity for the action of $\mathcal{G}(X)$ on $X$, and $|Z|$ is not a divisor of $|Y|$?
Below, we provide a family of examples that answer positively to the previous question.
Example 64.
Let $p$ be a prime number, with $p>2$, and $X$ be the simple affine cycle set of size $p^{p}$ obtained by 57 with $\lambda=\mu=1$ and $c:=(0,\ldots,0)$. Since $\psi$ has $-(x-1)^{p}$ as characteristic polynomial, the Jordan canonical form of $\psi$ exists in $\mathbb{Z}/p\mathbb{Z}$ and is an upper triangular matrix. Let :=$\mathcal{B}:=\{e_{1},\ldots,e_{p}\}$ the basis of $X$ that realizes the required form. Then, the subspace $H$ generated by the set $\{e_{1},\ldots,e_{p-1}\}$ is $\psi$-invariant and has dimension $p-1$, moreover $X=H\oplus<e_{p}>$ and it correspond bijectively to the set $Y\times Z$ given by $Y:=\{0\}\oplus<e_{p}>$ and $Z:=H\oplus\{0\}$, where the correspondence is given by $h+te_{p}\mapsto((0,te_{p}),(h,0))$ for all $h\in H$, $t\in\mathbb{Z}/p\mathbb{Z}$. By 34, $X$ has a complete blocks system $\{X_{y}\}_{y\in Y}$, where $X_{y}:=\{(y,z)\mid z\in Z\}$, for $y\in Y$, has size $p^{p-1}$, and with $|Y|=p$.